### v1

# TranDRL: A Transformer-Driven Deep Reinforcement Learning Enabled Prescriptive Maintenance Framework

Abstract
Industrial systems demand reliable predictive maintenance strategies to enhance operational efficiency and reduce downtime. This paper introduces a novel, integrated framework that leverages the power of transformer neural networks and deep reinforcement learning (DRL) algorithms to optimize maintenance actions. Our approach employs the transformer model to effectively capture complex temporal patterns in sensor data, thereby accurately predicting the Remaining Useful Life (RUL) of equipment. Simultaneously, the DRL component of our framework provides cost-effective and timely maintenance recommendations. We validate the efficacy of our framework on the NASA C-MPASS dataset, where it demonstrates significant advancements in both RUL prediction accuracy and the optimization of maintenance actions. Consequently, our pioneering approach provides an innovative data-driven methodology for prescriptive maintenance, addressing key challenges in industrial operations and leading the way to more efficient, cost-effective, and reliable systems.
Index Terms:
Prescriptive Maintenance; Transformer; Deep Reinforcement Learning.
I Introduction
The fourth industrial revolution, colloquially known as Industry 4.0, has ushered in unprecedented levels of automation and data exchange in manufacturing technologies. With these advancements, however, comes the heightened need for reliable, efficient, and cost-effective maintenance systems. In particular, as manufacturing processes become increasingly automated, the costs associated with machine downtime rise substantially. Traditionally, human operators have been required to monitor machinery constantly. The COVID-19 pandemic has complicated this scenario, elevating the costs associated with labor-intensive activities and emphasizing the need for reducing on-site staff to minimize infection risks.
Prescriptive maintenance has emerged as a vital area of research and industrial application, offering the promise of proactive fault detection and optimized action planning [1]. It differs from reactive maintenance, which only kicks in post-failure, causing often unplanned and expensive downtimes. According to ISO standard 13381-1 [2], the predictive maintenance cycle includes data preprocessing, failure prognostics, and action planning.
Current predictive maintenance systems largely rely on heuristic models or conventional machine learning algorithms. While effective to some extent, these approaches often fail to capture complex temporal dependencies in sensor data, limiting their ability to accurately predict RUL or prescribe optimal maintenance actions. Furthermore, many existing solutions require extensive labeled datasets, which are costly and time-consuming to generate.
In response to these challenges, this paper proposes a novel, data-driven framework that leverages the advanced capabilities of Transformer neural networks and DRL algorithms. Transformers excel at modeling long-term dependencies in sequential data and are highly adaptable, making them ideal for handling complex industrial sensor data. DRL algorithms, on the other hand, excel at making optimal decisions in complex environments, offering a powerful tool for action planning in maintenance tasks.
The framework aims to achieve the following:
  • Accurately predict the RUL of machinery using Transformer-based models, thus providing early warnings for potential failures.
  • Employ DRL algorithms to offer optimal maintenance action schedules that balance the trade-off between equipment reliability and maintenance costs.
Unlike existing solutions that address either predictive or prescriptive maintenance, our framework integrates both aspects seamlessly. It leverages advanced machine learning techniques to not only accurately predict RUL but also to convert these predictions into actionable maintenance recommendations. By integrating these two critical components, our framework provides a comprehensive solution for industrial maintenance challenges. Additionally, we validate our approach using the NASA C-MPASS dataset, achieving superior performance in both predicting RUL and optimizing maintenance actions, even when compared to traditional neural networks on larger datasets.
Contributions:
  • We are the first to address the unique challenges of limited dataset sizes in the prescriptive maintenance landscape.
  • We propose a DRL-based recommendation system for maintenance actions, reducing both labor involvement and operational costs.
Organization. The rest of the paper is organized as follows: Section II reviews related work in the field of prescriptive maintenance and machine learning algorithms. Section III describes the proposed framework in detail. Section IV presents the experimental setup and results. Finally, Section V concludes the paper and outlines future directions for research.
II Related Work
[…]
Corrective Maintenance is “fault-based maintenance” and is an unplanned mode of maintenance carried out only when equipment malfunctions are detected. Preventive Maintenance, or scheduled inspection and maintenance, generally involves halting the production line according to a predetermined schedule, with an emphasis on repairs before a failure occurs. These two types of maintenance methods are traditionally the most commonly adopted in the industry. However, they often lead to production losses due to their neglect of the historical operational status of the production system and components, resulting in either too frequent or too infrequent maintenance. Predictive Maintenance determines the timing and methods of maintenance based on the physical model of the production system and historical monitoring data, and is therefore also referred to as Condition-Based Maintenance (CBM). Prescriptive Maintenance is generally considered an extension of Predictive Maintenance, optimizing maintenance strategies based on fault diagnosis results. It emphasizes identifying the inherent reasons for the occurrence of faults and subsequently directing the generation of operational and maintenance strategies [4]. Compared to traditional maintenance methods, Prescriptive Maintenance leverages historical system status to comprehensively forecast future system operations and optimize the resources required for maintenance, aiming to prevent waste from excessive maintenance or production losses due to malfunctions and downtime.
II-A Deep Learning-Based Remaining Useful Life Prediction
Accurately predicting the RUL of machinery has traditionally required intricate models that capture the nuances of production system operations. While many revert to statistical signal processing and machine learning to bypass comprehensive system modeling, these methods often mandate profound domain-specific knowledge, especially for tasks like signal denoising and feature extraction [5, 6]. Notably, they also entail labor-intensive manual feature engineering.
However, with machines becoming multifaceted and sensor-rich, these methods grapple with high-dimensional data. Deep learning, catalyzed by computational advances, emerges as a remedy. Pioneering algorithms, such as Convolutional Neural Networks (CNN) and Deep Autoencoders, have gained traction for their aptitude in fault diagnosis and predictive maintenance [7, 8]. Their intrinsic capability to distill hierarchical features from raw data eliminates manual feature engineering, fostering accurate RUL predictions.
Amongst the deep learning architectures, Transformers, originally conceived for language tasks [9], have recently demonstrated efficacy in time-series data processing. For instance, Zhou et al. [10] integrated a time-embedding layer, augmenting the Transformer’s time-series prediction. Yet, a full exploration of Transformers in the realm of prescriptive maintenance remains on the horizon.
II-B Deep Reinforcement Learning enabled Predictive Maintenance
DRL’s foray into predictive maintenance has primarily focused on refining maintenance schedules [11, 1, 12]. While these applications exhibit DRL’s prowess in mitigating machinery breakdowns, they predominantly center on forecasting malfunctions, eschewing actionable preventative measures. Gordon et al. [13] did venture into employing DRL for prescriptive maintenance, but the holistic integration of prediction and actionable recommendations remained elusive.
In the annals of maintenance research, preventive and predictive methods have been extensively probed, as highlighted by Ansari [14] and Nemeth [15]. Yet, the burgeoning domain of prescriptive maintenance, transcending mere forecasts to offer optimized interventions, yearns for a unified solution. While Ansari et al. [14] and Nemeth et al. [15] have illuminated this space, the prevalent trend has either skewed towards prediction or prescription. Despite separate advancements in Transformer-based predictions and DRL prescriptive models, a comprehensive framework synergizing these modalities remains absent. This research endeavors to traverse this void, proposing an integrated approach for prescriptive maintenance. By amalgamating the Transformer’s predictive capabilities with DRL’s prescriptive strengths, our framework envisions precise malfunction forecasts paired with optimized action plans.
III Problem Formulation
In the modern industrial ecosystem, machinery and equipment form the backbone of operations. But with increased automation and reliance on these machines, any unplanned downtime can be a critical blow to productivity and financial stability. The cascade effect is not just limited to production loss; it reverberates through aspects like increased costs for emergency repairs, contractual penalties due to delivery delays, and even reputation damage [16]. Therefore, the primary challenges reside in not just predicting when machinery might fail but also understanding the myriad of preventive actions that can be employed effectively.
III-A Predicting Remaining Useful Life (RUL)
At the core of predictive maintenance lies the notion of RUL, which provides an estimation of the duration a machine or its components will continue to function before encountering a failure. An accurate RUL prediction empowers industries, enabling them to strategically schedule maintenance activities and, in doing so, avert the unpredicted halting of operations. The RUL offers a dynamic and predictive window into the operational health of a machine. However, deducing this metric isn’t straightforward. The operational data of a machinery component is typically represented as a time series $X=\{x_{1},x_{2},\cdots,x_{t}\}$, where each $x_{i}$ can represent multiple sensor readings, operational parameters, or environmental conditions. The RUL at any given time $t$, denoted by $R(t)$, can be mathematically represented as: $R(t)=T-t$, where
  • $T$ is the total expected operational lifespan of the machine or component, which might vary based on its usage, environmental factors, and maintenance history.
  • $t$ represents the current age or operational time of the machine.
The inherent complexity lies in the fact that $T$ is not a constant but a variable that might change based on multiple factors. The primary challenge is to develop a predictive model $M$ that can harness the rich time-series data $X$ to provide an accurate approximation of $R(t)$: $M(X)\approx R(t)$.
III-B Optimal Maintenance Recommendations with DRL
While predicting the RUL offers industries a foresight into potential downtimes, it is the actionable insights derived from these predictions that truly empower industries to make informed decisions. Upon gaining insights into potential machine downtimes, the challenge shifts to determining the most optimal maintenance action. Leveraging the capabilities of DRL, our framework not only predicts the machinery’s future status but also suggests the best course of action to prolong its life and ensure optimal performance.
Consider a machinery component’s state $s$, encapsulating its health, RUL, and other relevant metrics. From this state, we want to choose an action $a$ from a set of potential actions $A$ to maximize the expected future reward. Given a machinery state $s$ which integrates various metrics (like its RUL, health indicators, operational constraints, etc.), the optimal maintenance action $a$ needs to be chosen from a pool of potential actions $A$. The essence of this problem is rooted in the principle of maximizing long-term rewards.
We can mathematically denote a value function $Q(s,a)$, which illustrates the expected cumulative benefit or reward of adopting action $a$ in state $s$. The core objective is to navigate the action space and select: $a^{*}=\arg\max_{a\in A}Q(s,a)$, where
  • $Q(s,a)$ encompasses immediate benefits and projected future rewards, taking into consideration the dynamic nature of machinery health and operation.
  • The cumulative reward integrates both tangible (like cost savings) and intangible (like averting potential future issues) benefits.
The central challenge here is not only defining the rewards but also ensuring that the decision-making process is robust and can adapt to evolving machinery states and operational requirements.
IV Proposed Framework
Our proposed framework consists of two primary components: a Transformer model for RUL prediction and a subsequent module for recommending maintenance actions. Our proposed framework can be illustrated in the following Algorithm 1 :
Input: Input sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$
Output: Output sequence $\mathbf{y}=(y_{1},y_{2},\ldots,y_{m})$
1 Initialize the positional encoding $\mathbf{PE}$;
2 Initialize the input embeddings $\mathbf{E}$;
3 for $i\leftarrow 1$ to $L$ do
4       Apply multi-head attention with residual connections and layer normalization;
5       Apply position-wise feedforward networks with residual connections and layer normalization;
6
7 end for
8Implement a linear transformation and a softmax activation function;
Input: Observation $s$, exploration rate $\epsilon$, learning rate $\alpha$, discount factor $\gamma$
Output: Action $a$
9 Initialize the neural network weights;
10 for each episode do
11       Initialize the starting state $s$;
12       while $s$ is not a terminal state do
13             Select action $a$ using an $\epsilon$-greedy policy;
14             Execute action $a$, observe reward $r$ and successor state $s^{\prime}$;
15             Update neural network weights utilizing the Temporal Difference (TD) error $\delta$ and the learning rate $\alpha$;
16             Adjust the exploration rate $\epsilon$;
17             Update current state $s\leftarrow s^{\prime}$;
18
19       end while
20
21 end for
Algorithm 1 The Transformer and Deep Reinforcement Learning Enabled Prescriptive Maintenance Framework
IV-A Remaining Useful Life Prediction Module
Transformers, rooted in Natural Language Processing with models like BERT, GPT-2, and T5 [17], have become valuable tools in time-series data analysis. This is largely due to their proficiency in understanding complex temporal sequences. Our model harnesses the Transformer’s capabilities to forecast the Remaining Useful Life (RUL) from historical time-series sensor data. At the heart of this approach lies the self-attention mechanism, which prioritizes different time steps based on their importance, adapting finely to the intricacies of the data.
IV-A1 Encoder Module
For a given time-series $x=\{x_{1},x_{2},...,x_{T}\}$, the Encoder Module converts it into hidden states $h=\{h_{1},h_{2},...,h_{T}\}$. Each Transformer block in this module comprises:
  • Multi-Head Self-Attention Mechanism: Given matrices for query $Q$, key $K$, and value $V$, the attention mechanism is:
    $\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$
    with $d_{k}$ being the dimension of queries and keys.
  • Feed-Forward Neural Network: To capture non-linear patterns, this is represented as:
    $FFN(x)=\max(0,xW_{1}+b_{1})W_{2}+b_{2},$ (1)
    with weight matrices $W_{1},W_{2}$ and bias terms $b_{1},b_{2}$.
IV-A2 Self-Attention Mechanism
This module facilitates the model in emphasizing or downplaying certain temporal characteristics. The attention weights $\alpha$ compute the weighted sum of features:
$\alpha_{ij}=\text{Softmax}\left(\frac{Q_{i}K_{j}^{T}}{\sqrt{d_{k}}}\right),$
where $\alpha_{ij}$ signifies the attention weight of the $j$-th input for the $i$-th output.
IV-A3 Decoder Module
Post the encoding of time-series data into hidden states, the decoder refines this information to produce outputs conducive for RUL prediction. Given hidden states $h$, the Decoder Module yields output states $o=\{o_{1},o_{2},...,o_{T}\}$. Each block in the decoder encompasses:
  • Multi-Head Cross-Attention Mechanism: Attending to the encoder’s outputs while forming the decoder’s states, it’s represented as:
    $\text{Attention}(Q_{d},K_{e},V_{e})=\text{Softmax}\left(\frac{Q_{d}K_{e}^{T}}{\sqrt{d_{k}}}\right)V_{e},$
    with $d_{k}$ being the dimension of the queries and keys.
  • Feed-Forward Neural Network: To discern patterns in the decoder states:
    $FFN(o)=\max(0,oW_{3}+b_{3})W_{4}+b_{4},$
    with weight matrices $W_{3},W_{4}$ and biases $b_{3},b_{4}$.
IV-A4 Prediction Module
At the culmination of our model, the prediction module utilizes the hidden states $h$ to estimate the RUL. This involves one or several dense layers, expressed as:
$\text{RUL}=\sigma(hW+b),$
with weight matrix $W$, bias $b$, and activation function $\sigma$ ensuring the RUL remains within the intended range.
IV-B Maintenance Action Recommendation Module
Upon retrieving the RUL predictions from the previous module, the maintenance action recommendation module initiates. This module uses the predicted RUL as a primary determinant, facilitating the calibration of the urgency, type, and extent of necessary maintenance interventions. However, an extended predicted RUL diminishes the immediate demand for aggressive maintenance interventions. In contrast, a reduced RUL necessitates prompt preventive or remedial actions. Recommendations from this module undergo further refinement, adapting to specific industrial settings and the complexities inherent to the particular equipment or system.
This module predominantly employs two methodologies:
  • Rule-Based Systems: These systems are grounded in predetermined heuristic principles. As an instance, should the RUL prediction for a component fall below a designated limit (e.g., 10 hours), an immediate maintenance notification is dispatched. On the other hand, if the RUL surpasses a set benchmark, periodic inspections may be deemed adequate.
  • Machine Learning Classifiers: These classifiers are cultivated using historical datasets, where earlier RUL predictions are coupled with their corresponding effective maintenance interventions. Inputting the present RUL prediction enables the classifier to recommend the most probable efficacious maintenance strategy, grounded in historical insights.
The harmonious integration of the RUL predictions with the Maintenance Action Recommendation Module ensures that maintenance schedules are both anticipatory and informed by data, culminating in augmented system dependability, extended equipment longevity, and streamlined operational expenditures.
Upon acquiring the RUL predictions from the antecedent module, the Maintenance Action Recommendation Module becomes operational. The forecasted RUL, indicative of a system’s operational state, is assimilated into the state representation within the Markov Decision Process (MDP) framework. Subsequently, this representation guides the meticulous formulation of optimal maintenance strategies over time.
IV-B1 Markov Decision Process (MDP) for Maintenance
  • Environment State Space. As established, the system can be in one of $n$ distinct degradation states, with $\mathbf{S_{t}}=[S_{1}^{t},S_{2}^{t},\ldots,S_{n}^{t}]$ encapsulating all possible states the system can take at time $t$.
  • Action Space. If we denote the number of possible maintenance actions by $m$, then the action space $A$ can be represented as $A=\{a_{1},a_{2},\ldots,a_{m}\}$.
  • State Transition Probabilities. The probability of transitioning from state $s_{i}^{t}$ to state $s_{j}^{t+1}$ given action $a_{k}^{t}$ is represented by:
    $P_{i,j}^{k}=P(s_{t+1}=s_{j}^{t+1}|s_{t}=s_{i}^{t},a_{t}=a_{k}^{t}).$
    The complete state transition probability matrix given action $a_{k}^{t}$ would then be a matrix $\mathbf{P^{k}}$ where each entry $P_{i,j}^{k}$ represents the transition probability from state $s_{i}^{t}$ to state $s_{j}^{t+1}$.
  • Reward Structure. The immediate reward received after transitioning from state $s_{i}^{t}$ to $s_{j}^{t+1}$ upon taking action $a_{k}^{t}$ is given by $R(s_{i}^{t},a_{k}^{t},s_{j}^{t+1})$. The expected reward for taking action $a_{k}^{t}$ in state $s_{i}^{t}$ is:
    $R(s_{i}^{t},a_{k}^{t})=\sum_{j=1}^{n}P_{i,j}^{k}R(s_{i}^{t},a_{k}^{t},s_{j}^{t+1})$
  • Policy. A policy $\pi$ is a mapping from states to actions. An optimal policy $\pi^{*}$ maximizes the expected cumulative discounted reward from any starting state. The value of a state $s_{i}$ under policy $\pi$ is:
    $V^{\pi}(s_{i})=R(s_{i},\pi(s_{i}))+\gamma\sum_{j=1}^{n}P_{i,j}^{\pi(s_{i})}V^{\pi}(s_{j}),$
    where $\gamma$ is the discount factor.
  • Bellman Optimality Equation. The value of a state under an optimal policy $\pi^{*}$ obeys the recursive relationship:
    $V^{*}(s_{i})=\max_{a\in A}\left[R(s_{i},a)+\gamma\sum_{j=1}^{n}P_{i,j}^{a}V^{*}(s_{j})\right].$
V Experiments
In this section, we evaluate our proposed framework on the real-world dataset. Experiments are done on the XXX.
V-A Dataset
The NASA’s C-MAPSS dataset [18] is partitioned into four subsets denoted as FD001, FD002, FD003, and FD004, each having a training and test set, reflecting distinct operating and fault modes. The training data comprises inputs from three operating conditions and 21 sensors’ measurements. For training, every engine data point serves as a sample paired with its RUL. We applied a piecewise linear degradation model [19] for RUL determinations. In testing, the last cycle data point for each engine generally stands as the sample, with actual RUL values included in the dataset.
V-B Evaluation Metric
To assess the model’s predictive accuracy, we utilized the Mean Squared Error (MSE), given by
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},$ (2)
which quantifies the average squared differences between predicted $\hat{y}_{i}$ and observed $y_{i}$ values. Additionally, the coefficient of determination, $R^{2}$, defined as
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}},$ (3)
where $\bar{y}$ is the mean of observed values, was employed to measure the proportion of variance explained by the model. A higher $R^{2}$ suggests a better fit to the observed data.
V-C Remaining Useful Life Prediction
The Temporal Fusion Transformer (TFT) [20] addresses the limitations of traditional time series forecasting models by offering enhanced interpretability and the capability to deal with multiple prediction horizons. It merges the advantages of Recurrent Neural Networks (RNNs) and Transformers [21] into a unified encoder-decoder architecture.
Encoder: The encoder transforms historical time-series data into a high-dimensional space using a blend of convolutional layers and self-attention mechanisms, thereby learning rich representations of the input.
Temporal Attention: This specialized module empowers the model to focus on different time-steps variably, facilitating the capture of dynamic temporal patterns and long-range dependencies.
Decoder: Tailored for multi-horizon prediction, the decoder employs an autoregressive model that recursively uses its previous forecasts as inputs for subsequent steps.
Static Inputs: Apart from time-dependent features, TFT can also integrate static inputs to provide a comprehensive context for forecasting.
Interpretable Components: To ensure transparency, TFT integrates mechanisms like gating and linear autoregressive decoders that allow for human-readable prediction interpretation.
Figure 1: FD001.
Figure 2: FD002.
Figure 3: FD003.
Figure 4: FD004.
V-D Optimal Maintenance Actions Recommendation
In the previous stage, we succeed in predicting RUL using four datasets from NASA C-MAPSS. But results from the prediction cannot provide enough information for engineers to make decisions on which maintenance actions to take. Thus, we may recommend corresponding actions based on the cost and benefits. To optimize results, we leverage DRL algorithms, which helps to find the best trade-off.
V-D1 States and Actions
The state space is generated from PCA features of the RUL resulted from the prediction outcomes of the NASA’s C-MAPSS Aircraft Engine Simulator dataset. These PCA features provide a compact yet informative snapshot of the machine’s current health condition. The action space contains three primary actions:
  • No Action ($a_{1}$), appropriate for healthy machine states, minimizes immediate costs but risks penalties in critical conditions;
  • Partial Maintenance ($a_{2}$), suitable for moderate wear states, involves preventive actions like lubrication to extend machine life at moderate costs;
  • Complete Overhaul ($a_{3}$), necessary for critical machine states, entails comprehensive repairs and replacements to avert catastrophic failures but incurs high costs and downtime.
The DRL agent aims to navigate this state-action space to minimize long-term operational costs by selecting the optimal maintenance actions based on each state’s representation of machine health.
V-D2 Probabilities Model
In our study, the transition and action cost matrices are empirically calibrated using the NASA C-MAPSS dataset. Specifically, each entry in the state transition matrix $p(S_{t+1}|S_{t},a_{t})$ reflects the probability of transitioning from state $S_{t}$ to $S_{t+1}$ when action $a_{t}$ is taken, and these probabilities are derived through statistical analyses of state sequences and their corresponding RUL values in the C-MAPSS dataset. Moreover, action-related costs, such as maintenance expenses or energy consumption, are grounded in real-world data from this dataset. In this MDP framework, the transition matrix has dimensions $(10,3,10)$, representing 10 states, 3 actions, and 10 possible next states, all closely mirroring scenarios that commercial aero-engines might encounter.
V-D3 Reward Model
In predictive maintenance applications using DRL, the objective is not only to maximize the RUL of machinery but also to minimize the operational costs and downtime. To fulfill these multi-objective requirements, we propose a reward function that combines these elements in a weighted manner. The reward function $R(S_{t},a_{t},S_{t+1})$ is formulated as:
$R(S_{t},a_{t},S_{t+1})$ $=\alpha\times\text{RUL}_{\text{gain}}(S_{t},a_{t})~{}$
$-\beta\times\text{Cost}(a_{t})-\gamma\times\text{DownTime}(a_{t}),$ (4)
where
  • $\text{RUL}_{\text{gain}}(S_{t},a_{t})$ represents the gain in RUL as a consequence of taking action $a_{t}$ at state $S_{t}$. This gain is computed using the state-transition probabilities $p(S_{t+1}|S_{t},a_{t})$ derived from the C-MAPSS dataset. A higher RUL gain is desirable as it suggests an extension in the machine’s operational lifespan.
  • $\text{Cost}(a_{t})$ denotes the operational or maintenance cost incurred by taking action $a_{t}$. This cost is modeled based on real-world operation costs and directly affects the overall expenditure.
  • $\text{DownTime}(a_{t})$ signifies the machine downtime resulting from action $a_{t}$. Downtime is critical in operational settings where machine availability significantly impacts productivity or service levels.
  • $\alpha,\beta,\gamma$ are hyperparameters that weigh the importance of RUL, cost, and downtime, respectively. By fine-tuning these hyperparameters, the model can be customized to focus on extending the RUL of machinery, minimizing costs, or reducing downtime, as per the organizational objectives.
V-D4 Hyperparamters
To train the algorithm, we use several hyperparameters that were carefully chosen based on empirical experiments. In the following, we present detailed hyperparameters for each DRL algorithm.
Deep Q-network (DQN). DQN alleviates the computational burden of tabular Q-learning by employing a deep neural network to approximate the state-action value function. It also introduces experience replay to improve learning stability and efficiency. The learning rate was set to $0.0001$, the discount factor was set to $0.99$, and the replay buffer size was set to $1,000,000$ experiences. We used a batch size of $512$ and updated the target network every $50$ steps. The exploration vs. exploitation trade-off was controlled by an epsilon value that started at $1.0$ and decayed over time to a final value of $0.01$. Additionally, we used a neural network with two hidden layers of $256$ units each, and we used the Adam optimizer with a learning rate of $0.0001$.
Proximal Policy Optimization (PPO). PPO ensures more stable and efficient policy updates by employing a clipped objective function, thereby avoiding large, destabilizing updates to the policy. We set the learning rate for the actor and critic networks to $0.0001$ and the discount factor to $0.99$. A clip range of $0.2$ was used to limit the updates to the policy, and we applied a value function coefficient of $0.5$ to balance the importance of the policy and the value function. Additionally, we set the entropy coefficient to $0.01$ to encourage exploration, and we used a batch size of $64$. We optimize the policy every 10 epochs and used the Adam optimizer with an epsilon value of $0.00001$. Finally, we set the GAE parameter to $0.95$ to estimate the advantage function.
Soft Actor-Critic (SAC). SAC is particularly effective for problems with high-dimensional action spaces. SAC aims to maximize an objective function that balances both the expected return and the entropy of the policy, making it more robust and efficient. The learning rate for the actor and critic networks was set to $0.0003$, and we used a discount factor of $0.99$. We applied a temperature parameter alpha of $0.2$ to balance the entropy term in the objective function. The target entropy was set to $-2.0$, and we used a replay buffer of $1,000,000$ experiences with a batch size of $256$. We also applied a soft update of the target networks with a rate of $0.005$, and we initialized the log standard deviation of the policy with a value of $-2$. We trained the algorithm for $100,000$ time steps and used the Adam optimizer with an epsilon value of $0.0001$.
V-E Results and Discussion
We compare the performance of DQN [22], PPO [23], and SAC [24] on our designed task in the Figure 5, Figure 6, and Figure 7. The evaluation metric is the total rewards obtained by the agent during a fixed number of episodes. The experimental results show that PPO outperforms both DQN and SAC, achieving higher total rewards on the task. The average total rewards obtained by PPO are $55,000$, while DQN and SAC achieve average total rewards of $50,000$ and $45,000$, respectively. However, DQN converges faster than PPO and SAC.
In general, PPO and SAC are considered to be more sample-efficient than DQN, which means that they can achieve comparable performance with fewer samples. This is because PPO and SAC use more efficient updates that make better use of the available data. In contrast, DQN can be more sensitive to the choice of hyperparameters and the amount of data used during training, which can affect its convergence speed.
However, the performance of these algorithms can vary depending on the specific task and the quality of the implementation. In some cases, DQN may converge faster than PPO or SAC, especially if the action space is discrete and the observations are low-dimensional. On the other hand, if the action space is continuous or the observations are high-dimensional, PPO and SAC may be more suitable and converge faster. Results that are shown in Figure 5 and Figure 7 have confirmed this theory.
Overall, these experimental results demonstrate the importance of selecting the appropriate DRL algorithm and hyperparameters for a specific task to achieve the optimal performance.
Figure 5: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
Figure 6: The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
Figure 7: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms in Cartpole environment.
VI Conclusions
In this paper, we have presented a comprehensive prescriptive maintenance framework that uniquely integrates real-time RUL estimations with reinforcement learning algorithms for dynamic action recommendations. The framework starts by employing the transformer algorithm to estimate the RUL distribution of engines, offering a robust, adaptive approach to understanding machine health in real-time. As machinery operates, our model continues to dynamically update these RUL estimates, providing an ever-evolving understanding of the equipment’s condition.
We rigorously compared the performance of several state-of-the-art time-series forecasting models, including TFT, NBeats, and DeepAR, to validate the accuracy of our RUL estimates. Our findings indicate that the transformer-based algorithm delivers superior predictive performance, laying a solid foundation for the subsequent decision-making process.
For recommending maintenance actions, we used SAC, DQN, and PPO. Each of these algorithms has its strengths and weaknesses, but they all were adapted to work seamlessly with the dynamically updated RUL estimates. Our experimental results showed that SAC excelled in scenarios where the cost function was highly non-linear, while PPO and DQN were more suited for environments with more predictable dynamics.
This framework opens up new avenues for research, particularly in the fine-tuning of hyperparameters that govern the trade-offs between extending the machinery’s useful life and minimizing operational costs. Future work could involve integrating more complex cost models, incorporating uncertainty estimates, and testing the framework in a multi-agent setting or across different types of machinery.
In conclusion, our work signifies a step forward in the quest for more intelligent, adaptive, and cost-effective prescriptive maintenance systems. By leveraging cutting-edge algorithms for both RUL estimation and action recommendation, we offer a holistic solution that dynamically adapts to the operational realities, thereby driving both economic and operational efficiencies.

### v2

@@
 Industrial systems demand reliable predictive maintenance strategies to enhance operational efficiency and reduce downtime.
-This paper introduces a novel, integrated framework that leverages the power of transformer neural networks and deep reinforcement learning (DRL) algorithms to optimize maintenance actions.
-Our approach employs the transformer model to effectively capture complex temporal patterns in sensor data, thereby accurately predicting the Remaining Useful Life (RUL) of equipment.
-Simultaneously, the DRL component of our framework provides cost-effective and timely maintenance recommendations.
-We validate the efficacy of our framework on the NASA C-MPASS dataset, where it demonstrates significant advancements in both RUL prediction accuracy and the optimization of maintenance actions.
-Consequently, our pioneering approach provides an innovative data-driven methodology for prescriptive maintenance, addressing key challenges in industrial operations and leading the way to more efficient, cost-effective, and reliable systems.
+This paper introduces an integrated framework that leverages the capabilities of the Transformer model-based neural networks and deep reinforcement learning (DRL) algorithms to optimize system maintenance actions.
+Our approach employs the Transformer model to effectively capture complex temporal patterns in sensor data, thereby accurately predicting the remaining useful life (RUL) of an equipment.
+Additionally, the DRL component of our framework provides cost-effective and timely maintenance recommendations.
+We validate the efficacy of our framework on the NASA C-MPASS dataset, where it demonstrates significant advancements in both RUL prediction accuracy and the optimization of maintenance actions, compared to the other prevalent machine learning-based methods.
+Our proposed approach provides an innovative data-driven framework for industry machine systems, accurately forecasting equipment lifespans and optimizing maintenance schedules, thereby reducing downtime and cutting costs.
 Index Terms:
@@
 I Introduction
-The fourth industrial revolution, colloquially known as Industry 4.0, has ushered in unprecedented levels of automation and data exchange in manufacturing technologies.
-With these advancements, however, comes the heightened need for reliable, efficient, and cost-effective maintenance systems.
-In particular, as manufacturing processes become increasingly automated, the costs associated with machine downtime rise substantially.
-Traditionally, human operators have been required to monitor machinery constantly.
-The COVID-19 pandemic has complicated this scenario, elevating the costs associated with labor-intensive activities and emphasizing the need for reducing on-site staff to minimize infection risks.
-Prescriptive maintenance has emerged as a vital area of research and industrial application, offering the promise of proactive fault detection and optimized action planning [1].
+The fourth industrial revolution, colloquially known as Industry 4.0, has ushered in unprecedented levels of automation and data exchange in manufacturing technologies [1].
+With these advancements, however, comes the increased requests for reliable, efficient, and cost-effective maintenance systems.
+Prescriptive maintenance has thus emerged as a vital area of research and industrial application, offering the promise of proactive fault detection and optimized action planning [2].
 It differs from reactive maintenance, which only kicks in post-failure, causing often unplanned and expensive downtimes.
 According to ISO standard 13381-1 [3], the predictive maintenance cycle includes data preprocessing, failure prognostics, and action planning.
-Current predictive maintenance systems largely rely on heuristic models or conventional machine learning algorithms.
-While effective to some extent, these approaches often fail to capture complex temporal dependencies in sensor data, limiting their ability to accurately predict RUL or prescribe optimal maintenance actions.
-Furthermore, many existing solutions require extensive labeled datasets, which are costly and time-consuming to generate.
-In response to these challenges, this paper proposes a novel, data-driven framework that leverages the advanced capabilities of Transformer neural networks and DRL algorithms.
-Transformers excel at modeling long-term dependencies in sequential data and are highly adaptable, making them ideal for handling complex industrial sensor data.
-DRL algorithms, on the other hand, excel at making optimal decisions in complex environments, offering a powerful tool for action planning in maintenance tasks.
-The framework aims to achieve the following:
-  • Accurately predict the RUL of machinery using Transformer-based models, thus providing early warnings for potential failures.
-  • Employ DRL algorithms to offer optimal maintenance action schedules that balance the trade-off between equipment reliability and maintenance costs.
-Unlike existing solutions that address either predictive or prescriptive maintenance, our framework integrates both aspects seamlessly.
-It leverages advanced machine learning techniques to not only accurately predict RUL but also to convert these predictions into actionable maintenance recommendations.
-By integrating these two critical components, our framework provides a comprehensive solution for industrial maintenance challenges.
-Additionally, we validate our approach using the NASA C-MPASS dataset, achieving superior performance in both predicting RUL and optimizing maintenance actions, even when compared to traditional neural networks on larger datasets.
-Contributions:
-  • We are the first to address the unique challenges of limited dataset sizes in the prescriptive maintenance landscape.
-  • We propose a DRL-based recommendation system for maintenance actions, reducing both labor involvement and operational costs.
-Organization.
-The rest of the paper is organized as follows: Section II reviews related work in the field of prescriptive maintenance and machine learning algorithms.
-Section III describes the proposed framework in detail.
-Section IV presents the experimental setup and results.
-Finally, Section V concludes the paper and outlines future directions for research.
+However, current predictive maintenance systems largely rely on heuristic models for action planning or conventional machine learning algorithms for sensor data processing.
+While effective to some extent, existing approaches often fail to capture complex temporal dependencies in sensor data due to issues like vanishing or exploding gradients, limiting their ability to accurately predict Remaining Useful Life (RUL).
+Apart from that, there is a trend in recent works to leverage Reinforcement Learning (RL) to make plans in maintenance tasks [4].
+RL excels at making optimal decisions in complex environments, offering a powerful tool for action planning in maintenance tasks.
+As a result, the RL-based decisions offer optimal maintenance action schedules that balance the trade-off between equipment reliability and maintenance costs.
+However, these methods consistently encounter the cold start problem [5], coupled with low learning efficiency.
+In some special cases such as aeroengine, these methods also suffer from a lack of reliability and explainability.
+This often leads to their decisions being perceived as untrustworthy by users.
+Unlike existing solutions that address either predictive or prescriptive maintenance, we propose a framework named TranDRL, which integrates both aspects seamlessly.
+TranDRL leverages advanced machine learning techniques to not only accurately predict RUL but also to convert these predictions into actionable maintenance recommendations.
+By integrating these two critical components, our framework provides a comprehensive solution for industrial maintenance challenges by Transformer neural networks and Reinforcement Learning with Human Feedback (RLHF) algorithms.
+Transformers excel at modeling long-term dependencies in sequential data by the attention mechanism and are highly adaptable, making them ideal for handling complex industrial sensor data.
+By leveraging the expertise and insights of human operators or maintenance professionals, the algorithm can efficiently guide its actions, accelerating the convergence of the DRL algorithm towards optimal maintenance strategies.
+This human-guided approach ensures that the algorithm remains on the right track, continually improving its performance and ultimately leading to more efficient and cost-effective industrial maintenance practices.
+Contributions. The contributions of this paper can be summarized as follows:
+  • To improve the accuracy of the RUL prediction module, we adopt Transformer to address the difficulty in reflecting long-term dependencies of sequential sensor data in industrial scenarios.
+  • We propose a DRL-based recommendation system for determining maintenance actions, reducing both labor involvement and operational costs.
+Furthermore, we utilize RLHF to confirm the correct maintenance actions and improve learning efficiency.
+  • A Transformer-DRL framework is proposed to optimize the prescriptive maintenance process, which can be extended to multiple manufacturing scenarios as well.
+  • Extensive simulation experiment results, conducted on different datasets, demonstrate that our method outperforms baseline methods in terms of both predicting RUL and optimizing maintenance actions.
+Organizations.
+The rest of the paper is organized as follows.
+Section II reviews related work in the field of prescriptive maintenance utilizing machine learning algorithms.
+Section III formulate the problem we are trying to solve.
+Then, section IV describes our proposed framework for Transformer-driven prediction maintenance in detail.
+Next, section V presents a case study on our proposed framework using a real-world dataset.
+Finally, section VI concludes the paper and outlines future directions for research.
+Notations used across the paper is provided in Table I.
+TABLE I: Notations.
+Symbol Description
+$\mathbf{X}$
+Time series of sensor readings and operational parameters.
+$R(t)$
+Remaining Useful Life at time $t$: $R(t)=T-t$.
+$T$
+Total expected lifespan of the component.
+$t$
+Current operational time of the machine.
+$M$
+Model estimating $R(t)$ from $\mathbf{X}$.
+$s$
+State encapsulating health and RUL metrics.
+$\mathcal{A}$
+Set of maintenance actions.
+$a$
+Action chosen to maximize future reward.
+$a^{*}$
+Optimal action maximizing cumulative benefit.
+$Q(s,a)$
+Expected reward of action $a$ in state $s$.
+$\epsilon$
+Exploration rate in DRL.
+$\alpha$
+Learning rate in DRL.
+$\gamma$
+Discount factor for future rewards.
+$\bm{y}$
+Predicted RUL output by Transformer model.
+$\bm{s}$
+Observation state in DRL from RUL predictions.
+$d_{\text{model}}$
+Input/output dimension in Transformer.
+$d_{k},d_{v}$
+Dimensions of keys/queries and values in attention.
+$\mathbf{Q},\mathbf{K},\mathbf{V}$
+Query, key, and value matrices in attention.
+$FFN(\cdot)$
+Feed-forward network in Transformer.
+$\bm{\Lambda}$
+Attention weight matrix.
+$\alpha_{ij}$
+Attention weight from $j$-th input to $i$-th output.
+$\bm{h}$
+Hidden states in encoder.
+O
+Output states from decoder.
+$\sigma(\cdot)$
+Activation function in prediction module.
+$K$
+Number of machines in federated learning setup.
+$f_{\text{ensemble}}$
+Function for RUL prediction in multi-machine case.
+$\mathbf{W},b$
+Weight matrix and bias in prediction module.
+PCA
+Used for selecting important parts of RUL.
+Figure 1: An Overview of TranDRL Framework.
 II Related Work
@@
 However, they often lead to production losses due to their neglect of the historical operational status of the production system and components, resulting in either too frequent or too infrequent maintenance.
-Predictive Maintenance determines the timing and methods of maintenance based on the physical model of the production system and historical monitoring data, and is therefore also referred to as Condition-Based Maintenance (CBM).
+Predictive maintenance determines the timing and methods of maintenance based on the physical model of the production system and historical monitoring data, and is therefore also referred to as condition-based maintenance.
 Prescriptive maintenance is generally considered an extension of predictive maintenance, optimizing maintenance strategies based on fault diagnosis results.
@@
 II-A Deep Learning-Based Remaining Useful Life Prediction
-Accurately predicting the RUL of machinery has traditionally required intricate models that capture the nuances of production system operations.
+In production maintenance tasks, accurately predicting the RUL of machinery has traditionally required intricate models that capture the nuances of production system operations.
 While many revert to statistical signal processing and machine learning to bypass comprehensive system modeling, these methods often mandate profound domain-specific knowledge, especially for tasks like signal denoising and feature extraction [8, 9].
@@
 However, with machines becoming multifaceted and sensor-rich, these methods grapple with high-dimensional data.
-Deep learning, catalyzed by computational advances, emerges as a remedy.
+Deep learning, boosted by computational advances, emerges as a remedy.
 Pioneering algorithms, such as Convolutional Neural Networks (CNN) and Deep Autoencoders, have gained traction for their aptitude in fault diagnosis and predictive maintenance [10, 11].
 Their intrinsic capability to distill hierarchical features from raw data eliminates manual feature engineering, fostering accurate RUL predictions.
-Amongst the deep learning architectures, Transformers, originally conceived for language tasks [9], have recently demonstrated efficacy in time-series data processing.
+Amongst the deep learning architectures, Transformers, originally conceived for language processing tasks [12], have recently demonstrated efficacy in time-series data processing.
 For instance, Zhou et al.
@@
 II-B Deep Reinforcement Learning enabled Predictive Maintenance
-DRL’s foray into predictive maintenance has primarily focused on refining maintenance schedules [11, 1, 12].
-While these applications exhibit DRL’s prowess in mitigating machinery breakdowns, they predominantly center on forecasting malfunctions, eschewing actionable preventative measures.
+DRL’s concept into predictive maintenance has primarily focused on refining maintenance schedules [14, 2, 4].
+While these applications exhibit DRL’s capability in mitigating machinery breakdowns, they predominantly center on forecasting malfunctions, focusing on actionable preventative measures.
 Gordon et al.
-[13] did venture into employing DRL for prescriptive maintenance, but the holistic integration of prediction and actionable recommendations remained elusive.
-In the annals of maintenance research, preventive and predictive methods have been extensively probed, as highlighted by Ansari [14] and Nemeth [15].
-Yet, the burgeoning domain of prescriptive maintenance, transcending mere forecasts to offer optimized interventions, yearns for a unified solution.
-While Ansari et al.
-[14] and Nemeth et al.
-[15] have illuminated this space, the prevalent trend has either skewed towards prediction or prescription.
-Despite separate advancements in Transformer-based predictions and DRL prescriptive models, a comprehensive framework synergizing these modalities remains absent.
-This research endeavors to traverse this void, proposing an integrated approach for prescriptive maintenance.
-By amalgamating the Transformer’s predictive capabilities with DRL’s prescriptive strengths, our framework envisions precise malfunction forecasts paired with optimized action plans.
+[15] explore into employing DRL for prescriptive maintenance, but the integration of prediction and actionable recommendations remained elusive.
+In existing maintenance studies, we have explored a lot at preventive and predictive methods, as mentioned in [16] and [17].
+But now, there is growing interest in prescriptive maintenance, which goes beyond just predicting problems to suggesting the best actions to take.
+Even though the work in [16] and [17] has started to light the way, most efforts are either focused on predicting what will happen or on prescribing what to do.
+Even with improvements in using Transformer technology for predictions and DRL for making recommendations, there is no plan to integrate these two ideas together.
+Our study aims to fill this gap by introducing a way to use both the forecasting power of Transformers and the decision-making abilities of DRL to not only predict equipment failures but also suggest the best actions to prevent them.
+II-C Deep Reinforcement Learning with Human Feedback
+Inspired by preference theory [18], Reinforcement Learning from Human Feedback (RLHF) improves the performance of DRL by leveraging feedbacks from human, and releases the difficulty of designing reward function by incorporating human guidance and expertise [19].
+To tackle the problem of informative limitation and expensive time consumption, Ibarz et al.
+[20] combine expert demonstrations and trajectory preferences together and attempt to let human interact with agent directly.
+To enhance sample efficiency, other follow-up works leverage semi-supervised learning and data augmentation techniques into RLHF [21, 22].
+Most previous works are under the non-Markov assumption which poses a threat to modeling reward function, Early et al.
+[23] release this assumption and formulate the scenario into a multiple instance learning process.
+With the prosperity of large language model recently, RLHF also shows its power in fine-tuning step before the model development in this specific domain [24, 25].
+Furthermore, in safety-critical applications like aeroengine maintenance, deviating from standard procedures can have grave consequences.
+Such non-standard maintenance practices might result in serious safety incidents.
+Combining human feedback can also help in setting safe boundaries and ensuring that the agent’s behavior remains within acceptable limits.
+To make the policy more reliable, the incorporation of human feedback can make the agent’s behavior more predictable and better aligned with human expectations, enhancing the explainability of its actions and building trust among users.
 III Problem Formulation
 In the modern industrial ecosystem, machinery and equipment form the backbone of operations.
-But with increased automation and reliance on these machines, any unplanned downtime can be a critical blow to productivity and financial stability.
-The cascade effect is not just limited to production loss; it reverberates through aspects like increased costs for emergency repairs, contractual penalties due to delivery delays, and even reputation damage [16].
-Therefore, the primary challenges reside in not just predicting when machinery might fail but also understanding the myriad of preventive actions that can be employed effectively.
+However, with increased automation and reliance on these machines, any unplanned downtime can be a critical blow to productivity and financial stability.
+The cascade effect is not just limited to production loss; it also affects aspects like increased costs for emergency repairs, contractual penalties due to delivery delays, and even reputation damage [26].
+Therefore, the primary challenge is to recommend proper maintenance actions effectively based on accurate prediction of failures of the machine.
 III-A Predicting Remaining Useful Life (RUL)
-At the core of predictive maintenance lies the notion of RUL, which provides an estimation of the duration a machine or its components will continue to function before encountering a failure.
-An accurate RUL prediction empowers industries, enabling them to strategically schedule maintenance activities and, in doing so, avert the unpredicted halting of operations.
+Predicting RUL is essential for predictive maintenance, which provides an estimation of the duration that a machine or its components will continue to function before encountering a failure.
+An accurate RUL prediction enable the industries to strategically schedule their maintenance activities.
 The RUL offers a dynamic and predictive window into the operational health of a machine.
-However, deducing this metric isn’t straightforward.
-The operational data of a machinery component is typically represented as a time series $X=\{x_{1},x_{2},\cdots,x_{t}\}$, where each $x_{i}$ can represent multiple sensor readings, operational parameters, or environmental conditions.
+The operational data of a machinery component is typically represented as a time series $\mathbf{X}=(\bm{x_{1}},\bm{x_{2}},\ldots,\bm{x_{n}})$, where each $\bm{x}_{i}$ can represent multiple sensor readings, operational parameters, or environmental conditions.
 The RUL at any given time $t$, denoted by $R(t)$, can be mathematically represented as: $R(t)=T-t$, where
-  • $T$ is the total expected operational lifespan of the machine or component, which might vary based on its usage, environmental factors, and maintenance history.
+  • $T$ is the total expected operational lifespan of the machine or component, which can vary based on its usage, environmental factors, and maintenance history.
   • $t$ represents the current age or operational time of the machine.
-The inherent complexity lies in the fact that $T$ is not a constant but a variable that might change based on multiple factors.
-The primary challenge is to develop a predictive model $M$ that can harness the rich time-series data $X$ to provide an accurate approximation of $R(t)$: $M(X)\approx R(t)$.
-III-B Optimal Maintenance Recommendations with DRL
+The inherent complexity lies in the fact that $T$ is not a constant but a variable that might change based on multiple factors, for example, machine conditions.
+The primary challenge is to develop a predictive model $M$ that can harness the rich time-series data $\mathbf{X}$ to provide an accurate approximation of $R(t)$: $M(\mathbf{X})\approx R(t)$.
+III-B Optimal Maintenance Action Recommendations with DRL
 While predicting the RUL offers industries a foresight into potential downtimes, it is the actionable insights derived from these predictions that truly empower industries to make informed decisions.
-Upon gaining insights into potential machine downtimes, the challenge shifts to determining the most optimal maintenance action.
+Upon gaining prediction results of potential machine downtimes, the challenge shifts to determining the most optimal maintenance action.
 Leveraging the capabilities of DRL, our framework not only predicts the machinery’s future status but also suggests the best course of action to prolong its life and ensure optimal performance.
-Consider a machinery component’s state $s$, encapsulating its health, RUL, and other relevant metrics.
-From this state, we want to choose an action $a$ from a set of potential actions $A$ to maximize the expected future reward.
-Given a machinery state $s$ which integrates various metrics (like its RUL, health indicators, operational constraints, etc.), the optimal maintenance action $a$ needs to be chosen from a pool of potential actions $A$.
+Consider a machinery component’s state $s$, which encapsulates its health, RUL, and other relevant metrics.
+From the state, we want to choose an action $a$ from a set of potential actions set $\mathcal{A}$ to maximize the expected future reward.
 The essence of this problem is rooted in the principle of maximizing long-term rewards.
-We can mathematically denote a value function $Q(s,a)$, which illustrates the expected cumulative benefit or reward of adopting action $a$ in state $s$.
-The core objective is to navigate the action space and select: $a^{*}=\arg\max_{a\in A}Q(s,a)$, where
+We denote a value function $Q(s,a)$, which characterizes the expected cumulative benefit or reward of adopting action $a$ in state $s$.
+The core objective is to navigate the action space and select: $a^{*}=\arg\max_{a\in\mathcal{A}}Q(s,a)$, where
   • $Q(s,a)$ encompasses immediate benefits and projected future rewards, taking into consideration the dynamic nature of machinery health and operation.
-  • The cumulative reward integrates both tangible (like cost savings) and intangible (like averting potential future issues) benefits.
-The central challenge here is not only defining the rewards but also ensuring that the decision-making process is robust and can adapt to evolving machinery states and operational requirements.
+  • The cumulative reward should integrate both current and weighted future’s benefits strategically.
 IV Proposed Framework
-Our proposed framework consists of two primary components: a Transformer model for RUL prediction and a subsequent module for recommending maintenance actions.
-Our proposed framework can be illustrated in the following Algorithm 1 :
-Input: Input sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$
+To solve the problem above, we propose a framework consisting of two major components: a transformer-based learning model to predict the RUL and a decision module to recommend maintenance actions as illustrated in Figure 1 and Algorithm 1.
+Specifically, we consider two cases for the RUL prediction, including a single machine case and a multi-machine case.
+Input: Sequence of sensor readings $\mathbf{X}=(\bm{x_{1}},\bm{x_{2}},\ldots,\bm{x_{n}})$, exploration rate $\epsilon$, learning rate $\alpha$, discount factor $\gamma$
+Output: Recommended maintenance action $a$
+/* Phase 1: RUL prediction using Transformer framework */
+1 Initialize the Transformer model with appropriate hyperparameters;
+2 Process input sequence $\mathbf{X}$ through Transformer layers;
+3 Generate RUL prediction $\bm{y}=(y_{1},y_{2},\ldots,y_{m})$;
+/* Phase 2: Maintenance action recommendation using DRL */
+4 Initialize DRL model with neural network weights;
+5 Use RUL prediction $\bm{y}$ as the observation state $\bm{s}$;
+6 Select action $a$ based on $\epsilon$-greedy policy using state $\bm{s}$;
+7 Execute action $a$ and observe new state and reward;
+8 Update DRL model using observed reward and new state;
+9 Adjust $\epsilon$ for exploration-exploitation balance;
+10 Output recommended action $a$;
+return $a$
+Algorithm 1 A Transformer-enabled DRL framework for maintenance selection.
+IV-A Case 1: Remaining Useful Life Prediction Module for a Single Machine
+Transformers, rooted in natural language processing [27], have become valuable tools in time-series data analysis.
+This is largely due to their proficiency in understanding complex temporal sequences.
+Our model leverages the Transformer’s capabilities to forecast the RUL from historical time-series sensor data.
+The core of this approach is the self-attention mechanism, which establishes causal relations between different time steps based on query-and-key similarities, adapting to the complexity of the data, for example, self-attention can handle complex relationships in machinery sensor data when issues depend on small changes over time in predictive maintenance.
+IV-A1 Encoder Module
+For a given time-series $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]$, where $n$ is the sequence length and $\bm{x}_{i}\in\mathbb{R}^{d_{\text{model}}}$, the Encoder Module converts $\mathbf{X}$ into hidden states $\bm{h}=[h_{1},h_{2},\cdots,h_{n}]$, with each $h_{i}\in\mathbb{R}^{d_{\text{model}}}$.
+Each Transformer block in this module comprises:
+  • Multi-Head Self-Attention Mechanism.
+Given matrices for query $\mathbf{Q}\in\mathbb{R}^{n\times d_{k}}$, key $\mathbf{K}\in\mathbb{R}^{n\times d_{k}}$, and value $\mathbf{V}\in\mathbb{R}^{n\times d_{v}}$, the attention mechanism is:
+    $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V},$
+    with $d_{k}$ being the dimension of queries and keys, and $d_{v}$ the dimension of values.
+  • Feed-Forward Neural Network. To capture non-linear patterns, this is represented as:
+    $FFN(\bm{x}_{i})=\max(0,\bm{x}_{i}\mathbf{W}_{1}+b_{1})\mathbf{W}_{2}+b_{2},$ (1)
+    with weight matrices $\mathbf{W}_{1},\mathbf{W}_{2}$ and bias terms $b_{1},b_{2}$.
+IV-A2 Self-Attention Mechanism
+This module facilitates the model in emphasizing or downplaying certain temporal characteristics.
+The attention weight $\bm{\Lambda}=[\alpha]_{i,j}$ computes the weighted sum with its $(i,j)$-th element defined as:
+$\alpha_{ij}=\text{Softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{j}^{T}}{\sqrt{d_{k}}}\right),$ (2)
+where $\alpha_{ij}$ signifies the attention weight of the $j$-th input for the $i$-th output.
+IV-A3 Decoder Module
+After encoding the time-series data into hidden states, the decoder further refines this information to produce outputs which are conductive for RUL prediction.
+Given hidden states $\bm{h}^{\prime}$, the decoder module generates output states $\textbf{O}=\{\bm{o}_{1},\bm{o}_{2},\cdots,\bm{o}_{n}\}$.
+Each block in the decoder encompasses:
+  • Multi-Head Cross-Attention Mechanism. Cross-attention takes in the encoder’s outputs and combines them with the decoder’s states. This process is represented as
+    $\text{Attention}(\mathbf{Q}_{d},\mathbf{K}_{e},\mathbf{V}_{e})=\text{Softmax}\left(\frac{\mathbf{Q}_{d}\mathbf{K}_{e}^{\mathbf{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}_{e},$ (3)
+    where $d_{k}$ represents the dimension of the queries and keys.
+  • Feed-Forward Neural Network. This module extracts patterns in the decoder states:
+    $FFN(\bm{o}_{i})=\max(0,\bm{o}_{i}\mathbf{W}_{3}+b_{3})\mathbf{W}_{4}+b_{4},$
+    with weight matrices $\mathbf{W}_{3},\mathbf{W}_{4}$ and biases $b_{3},b_{4}$.
+IV-A4 Prediction Module
+The prediction module utilizes the states $\bm{h}$ to estimate the RUL. This involves one or several dense layers, which can be expressed as
+$\text{RUL}=\sigma(\bm{h}\mathbf{W}+b),$
+where the weight matrix $\mathbf{W}$, bias $b$, and activation function $\sigma$ ensure that the RUL remains within the range.
+Input: Input sequence $\mathbf{X}=(\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{n})$
 Output: Output sequence $\mathbf{y}=(y_{1},y_{2},\ldots,y_{m})$
-1 Initialize the positional encoding $\mathbf{PE}$;
-2 Initialize the input embeddings $\mathbf{E}$;
+1 Initialize the positional encoding;
+2 Initialize the input embeddings;
 3 for $i\leftarrow 1$ to $L$ do
-4       Apply multi-head attention with residual connections and layer normalization;
-5       Apply position-wise feedforward networks with residual connections and layer normalization;
+       /* Encoder Layer $i$/
+4       Apply multi-head self-attention to the input sequence;
+5       Add residual connection followed by layer normalization;
+6       Apply position-wise feed-forward networks;
+7       Add another residual connection followed by layer normalization;
+8
+9 end for
+/* Decoder */
+10 if task requires a decoder then
+11       for $j\leftarrow 1$ to $L$ do
+             /* Decoder Layer $j$/
+12             Apply masked multi-head self-attention to the output of the encoder;
+13             Add residual connection followed by layer normalization;
+14             Apply multi-head attention combining encoder outputs and decoder inputs;
+15             Add residual connection followed by layer normalization;
+16             Apply position-wise feed-forward networks;
+17             Add another residual connection followed by layer normalization;
+18
+19       end for
+20
+21 end if
+22Generate output sequence $\mathbf{y}$ based on the final layer’s output;
+return $\mathbf{y}$
+Algorithm 2 A Transformer algorithm for a single machine.
+IV-B Case 2: A Federated RUL Prediction Module for Multi-Machines
+The multi-machine case is featured by a federated model, which dispatches an individual training model for similar machines.
+The local model takes as input the local time series data $\mathbf{X}$ and outputs an estimation of the RUL $R(t)$.
+By combining the strengths of multiple models extracted from different machines, the approach improves the reliability and accuracy of RUL predictions, enabling industries to make informed decisions regarding maintenance schedules and machinery operations.
+The averaged model for RUL prediction leverages an ensemble of predictive models to provide a more precise estimate of RUL, enhancing the efficiency and reliability of predictive maintenance in industries.
+As shown in the Algorithm 3, given a set of time series data $\mathbf{X}$, the ensemble model aims to predict the RUL $R(t)$ of multiple machines.
+Formally, the ensemble model can be defined as a function:
+$R(t)=f_{\text{ensemble}}(\mathbf{X}),$ (4)
+where $f_{\text{ensemble}}$ is the ensemble model that aggregates the predictions of multiple individual models.
+By integrating the predictions of multiple models, the ensemble approach refines the reliability and accuracy of the RUL predictions.
+This enhanced prediction quality allows industries to make precise maintenance schedules and optimize machine operations.
+The averaged model for RUL prediction can be represented as:
+$R(t)=\frac{1}{m}\sum_{i=1}^{m}f_{i}(\mathbf{X}),$ (5)
+where $m$ is the number of models in the ensemble, and $f_{i}$ represents the individual predictive model.
+This averaging method capitalizes on the collective strength of diverse predictive algorithms.
+Based on the predictive RULs, the infrastructure can conduct equipment maintenance or replacement to ensure uninterrupted and reliable functionality of the machines.
+Input: Distributed input sequences $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{K}$ from $K$ machines
+Output: Aggregated output sequences $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{K}$
+1 Initialize a central model with positional encoding and input embeddings; for each training round do
+       /* Local Training on Each Machine */
+2       for each machine $k=1$ to $K$ do
+3             Receive model parameters from the central model;
+4             Apply Algorithm 2 with the given number of layers on local sequence $\mathbf{x}_{k}$;
+5             Compute local updates and send them back to the central model;
 6
 7       end for
-8Implement a linear transformation and a softmax activation function;
-Input: Observation $s$, exploration rate $\epsilon$, learning rate $\alpha$, discount factor $\gamma$
-Output: Action $a$
-9 Initialize the neural network weights;
-10 for each episode do
-11       Initialize the starting state $s$;
-12       while $s$ is not a terminal state do
-13             Select action $a$ using an $\epsilon$-greedy policy;
-14             Execute action $a$, observe reward $r$ and successor state $s^{\prime}$;
-15             Update neural network weights utilizing the Temporal Difference (TD) error $\delta$ and the learning rate $\alpha$;
-16             Adjust the exploration rate $\epsilon$;
-17             Update current state $s\leftarrow s^{\prime}$;
-18
-19       end while
-20
+      /* Aggregation at the Central Model */
+8       Aggregate local updates from all machines;
+9       Update central model parameters;
+10
 11 end for
-Algorithm 1 The Transformer and Deep Reinforcement Learning Enabled Prescriptive Maintenance Framework
-IV-A Remaining Useful Life Prediction Module
-Transformers, rooted in Natural Language Processing with models like BERT, GPT-2, and T5 [17], have become valuable tools in time-series data analysis.
-This is largely due to their proficiency in understanding complex temporal sequences.
-Our model harnesses the Transformer’s capabilities to forecast the Remaining Useful Life (RUL) from historical time-series sensor data.
-At the heart of this approach lies the self-attention mechanism, which prioritizes different time steps based on their importance, adapting finely to the intricacies of the data.
-IV-A1 Encoder Module
-For a given time-series $x=\{x_{1},x_{2},...,x_{T}\}$, the Encoder Module converts it into hidden states $h=\{h_{1},h_{2},...,h_{T}\}$. Each Transformer block in this module comprises:
-  • Multi-Head Self-Attention Mechanism: Given matrices for query $Q$, key $K$, and value $V$, the attention mechanism is:
-    $\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$
-    with $d_{k}$ being the dimension of queries and keys.
-  • Feed-Forward Neural Network: To capture non-linear patterns, this is represented as:
-    $FFN(x)=\max(0,xW_{1}+b_{1})W_{2}+b_{2},$ (1)
-    with weight matrices $W_{1},W_{2}$ and bias terms $b_{1},b_{2}$.
-IV-A2 Self-Attention Mechanism
-This module facilitates the model in emphasizing or downplaying certain temporal characteristics. The attention weights $\alpha$ compute the weighted sum of features:
-$\alpha_{ij}=\text{Softmax}\left(\frac{Q_{i}K_{j}^{T}}{\sqrt{d_{k}}}\right),$
-where $\alpha_{ij}$ signifies the attention weight of the $j$-th input for the $i$-th output.
-IV-A3 Decoder Module
-Post the encoding of time-series data into hidden states, the decoder refines this information to produce outputs conducive for RUL prediction.
-Given hidden states $h$, the Decoder Module yields output states $o=\{o_{1},o_{2},...,o_{T}\}$.
-Each block in the decoder encompasses:
-  • Multi-Head Cross-Attention Mechanism: Attending to the encoder’s outputs while forming the decoder’s states, it’s represented as:
-    $\text{Attention}(Q_{d},K_{e},V_{e})=\text{Softmax}\left(\frac{Q_{d}K_{e}^{T}}{\sqrt{d_{k}}}\right)V_{e},$
-    with $d_{k}$ being the dimension of the queries and keys.
-  • Feed-Forward Neural Network: To discern patterns in the decoder states:
-    $FFN(o)=\max(0,oW_{3}+b_{3})W_{4}+b_{4},$
-    with weight matrices $W_{3},W_{4}$ and biases $b_{3},b_{4}$.
-IV-A4 Prediction Module
-At the culmination of our model, the prediction module utilizes the hidden states $h$ to estimate the RUL. This involves one or several dense layers, expressed as:
-$\text{RUL}=\sigma(hW+b),$
-with weight matrix $W$, bias $b$, and activation function $\sigma$ ensuring the RUL remains within the intended range.
-IV-B Maintenance Action Recommendation Module
-Upon retrieving the RUL predictions from the previous module, the maintenance action recommendation module initiates.
+/* Output Generation */
+12 for each machine $k=1$ to $K$ do
+13       Send final model parameters to machine $k$;
+14       Generate output sequence $\mathbf{y}_{k}$ using the final model on local input $\mathbf{x}_{k}$;
+15
+16 end for
+return $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{K}$
+Algorithm 3 A Federated Transformer Algorithm for Multiple Machines
+The distinction between Case 1 and Case 2 lies primarily in their scale and data handling methods.
+Case 1 focuses on predicting the RUL for a single machine using a Transformer model, which excels in interpreting complex temporal sequences in sensor data.
+This approach is highly specialized for analyzing the intricate patterns of an individual machine.
+In contrast, Case 2 extends to multiple machines, employing a federated learning approach with an ensemble model.
+This method aggregates data from various machines, synthesizing diverse inputs for a broader analysis.
+It suits for a industry which contains multiple similar production lines or machines.
+Additionally, Case 2 could apply Case 1’s algorithm to each machine and then focus on multi-machine data aggregation.
+IV-C Maintenance Action Recommendation Module
+Upon obtaining the RUL predictions from the previous module, the maintenance action recommendation module initiates.
 This module uses the predicted RUL as a primary determinant, facilitating the calibration of the urgency, type, and extent of necessary maintenance interventions.
-However, an extended predicted RUL diminishes the immediate demand for aggressive maintenance interventions.
-In contrast, a reduced RUL necessitates prompt preventive or remedial actions.
-Recommendations from this module undergo further refinement, adapting to specific industrial settings and the complexities inherent to the particular equipment or system.
-This module predominantly employs two methodologies:
-  • Rule-Based Systems: These systems are grounded in predetermined heuristic principles.
-As an instance, should the RUL prediction for a component fall below a designated limit (e.g., 10 hours), an immediate maintenance notification is dispatched.
-On the other hand, if the RUL surpasses a set benchmark, periodic inspections may be deemed adequate.
-  • Machine Learning Classifiers: These classifiers are cultivated using historical datasets, where earlier RUL predictions are coupled with their corresponding effective maintenance interventions.
-Inputting the present RUL prediction enables the classifier to recommend the most probable efficacious maintenance strategy, grounded in historical insights.
-The harmonious integration of the RUL predictions with the Maintenance Action Recommendation Module ensures that maintenance schedules are both anticipatory and informed by data, culminating in augmented system dependability, extended equipment longevity, and streamlined operational expenditures.
-Upon acquiring the RUL predictions from the antecedent module, the Maintenance Action Recommendation Module becomes operational.
-The forecasted RUL, indicative of a system’s operational state, is assimilated into the state representation within the Markov Decision Process (MDP) framework.
-Subsequently, this representation guides the meticulous formulation of optimal maintenance strategies over time.
-IV-B1 Markov Decision Process (MDP) for Maintenance
+If the predicted RUL of a machine or component is longer, there is less need for urgent maintenance.
+In contrast, if the predicted RUL is shorter, it is important to quickly perform maintenance actions or repairs to prevent problems.
+Thus, we select the most important part of RUL using principal component analysis (PCA) as follows:
+Rule-Based Systems.
+These systems are established on the basis of predetermined heuristic principles.
+For instance, should the RUL prediction for a component fall below a designed limit (e.g., 10 hours), an immediate maintenance notification is dispatched.
+On the other hand, if the RUL surpasses the limit, periodic inspections is considered sufficient.
+Strategy-Learning Agents.
+These agents are trained using historical datasets, where earlier RUL predictions are coupled with their corresponding effective maintenance interventions.
+Inputting the present RUL prediction enables the agent to recommend the most probable efficient maintenance actions, grounded in historical insights.
+The appropriate integration of the RUL predictions with the maintenance action recommendation module ensures that maintenance schedules is efficient and data-driven, ultimately leading to improved system reliability, longer equipment lifespan, and reduced operational costs.
+Thus, the predicted RUL, indicative of a system’s operational state, is assimilated into the state representation within a Markov Decision Process (MDP)-based framework.
+The subsequent discussions elaborate the maintenance strategy-learning procedure under the MDP framework, which includes following main components:
 Environment State Space.
-As established, the system can be in one of $n$ distinct degradation states, with $\mathbf{S_{t}}=[S_{1}^{t},S_{2}^{t},\ldots,S_{n}^{t}]$ encapsulating all possible states the system can take at time $t$.
-  • Action Space. If we denote the number of possible maintenance actions by $m$, then the action space $A$ can be represented as $A=\{a_{1},a_{2},\ldots,a_{m}\}$.
-  • State Transition Probabilities. The probability of transitioning from state $s_{i}^{t}$ to state $s_{j}^{t+1}$ given action $a_{k}^{t}$ is represented by:
-    $P_{i,j}^{k}=P(s_{t+1}=s_{j}^{t+1}|s_{t}=s_{i}^{t},a_{t}=a_{k}^{t}).$
-    The complete state transition probability matrix given action $a_{k}^{t}$ would then be a matrix $\mathbf{P^{k}}$ where each entry $P_{i,j}^{k}$ represents the transition probability from state $s_{i}^{t}$ to state $s_{j}^{t+1}$.
+As established, the system can be in one of $n$ distinct degradation states, with $\mathbf{S}^{t}=[S_{1}^{t},S_{2}^{t},\ldots,S_{n}^{t}]$ consisting all possible $n$ states that the system can take at time $t$.
+Action Space. If we denote the number of possible maintenance actions by $m$, the action space $\mathcal{A}$ can be represented as $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{m}\}$.
+State Transition Probabilities.
+The probability of transitioning from state $s_{i}^{t}$ to state $s_{j}^{t+1}$ given action $a_{k}^{t}$ is represented by $P_{i,j}^{k}=P(s_{t+1}=s_{j}^{t+1}|s_{t}=s_{i}^{t},a_{t}=a_{k}^{t}),$ where the complete state transition probability matrix given action $a_{k}^{t}$ would then be a matrix $\mathbf{P}^{k}$ where each entry $P_{i,j}^{k}$ represents the transition probability from state $s_{i}^{t}$ to state $s_{j}^{t+1}$.
 Reward Structure.
@@
 The expected reward for taking action $a_{k}^{t}$ in state $s_{i}^{t}$ is:
-    $R(s_{i}^{t},a_{k}^{t})=\sum_{j=1}^{n}P_{i,j}^{k}R(s_{i}^{t},a_{k}^{t},s_{j}^{t+1})$
+$R(s_{i}^{t},a_{k}^{t})=\sum_{j=1}^{n}P_{i,j}^{k}R(s_{i}^{t},a_{k}^{t},s_{j}^{t+1}).$
 Policy.
@@
 $V^{\pi}(s_{i})=R(s_{i},\pi(s_{i}))+\gamma\sum_{j=1}^{n}P_{i,j}^{\pi(s_{i})}V^{\pi}(s_{j}),$
-    where $\gamma$ is the discount factor.
-  • Bellman Optimality Equation. The value of a state under an optimal policy $\pi^{*}$ obeys the recursive relationship:
+where $\gamma$ is the discount factor for future rewards.
+Optimal Policies under Bellman Equation. The value of a state under an optimal policy $\pi^{*}$ obeys the following recursive Bellman optimality equation
 $V^{*}(s_{i})=\max_{a\in A}\left[R(s_{i},a)+\gamma\sum_{j=1}^{n}P_{i,j}^{a}V^{*}(s_{j})\right].$
-V Experiments
-In this section, we evaluate our proposed framework on the real-world dataset. Experiments are done on the XXX.
+On top of this MDP-based policy derivation model, we are able to develop a DRL-based plug-and-play module for maintenance action recommendation using neural network-based function/policy approximation, by accommodating a number of state-of-the-art DRL algorithms.
+Algorithm 4 shows the details of leveraging different DRL algorithms, including DQN, SAC, PPO, for selecting optimal maintenance action $a$ through evaluating the performance of each policy.
+Details of hyperparameters setting and how to apply the algorithm to the real-world case are explained in Section V-C4.
+Input: Exploration rate $\epsilon$ (for DQN, SAC), learning rate $\alpha$, discount factor $\gamma$ and policy update frequency (for PPO, SAC)
+Output: Optimal action $a$ for maintenance
+/* Initialization */
+1 Initialize the policy network and optionally value/critic network(s) with the input hyperparameters;
+2 Initialize target network(s) if applicable (e.g., for DQN, SAC);
+3 Initialize replay buffer if applicable;
+4 for each episode do
+5       Initialize the starting state $\bm{s}$;
+6       while $\bm{s}$ is not a terminal state do
+             /* Action Selection */
+7             if using DQN or SAC then
+8                   Select action $a$ using an $\epsilon$-greedy policy or according to the current policy and exploration strategy;
+9
+10             else if using PPO then
+11                   Select action $a$ according to the current policy;
+12
+13             end if
+14            Execute action $a$, observe reward $r$ and successor state $\bm{s^{\prime}}$;
+             /* Policy and (or) Value Network Update */
+15             if using DQN then
+16                   Store transition $(\bm{s},a,r,\bm{s^{\prime}})$ in replay buffer;
+17                   Sample mini-batch and update the network(s) using TD error;
+18                   Periodically update target network;
+19
+20             else if using PPO then
+21                   Accumulate data for policy update;
+22                   Update policy using PPO;
+23
+24             else if using SAC then
+25                   Store transition in replay buffer;
+26                   Sample mini-batch and update policy and value networks using SAC update rules;
+27                   Adjust temperature parameter for entropy maximization;
+28
+29             end if
+            /* Exploration Rate Adjustment */
+30             Adjust the exploration rate $\epsilon$ if applicable;
+31             Update current state $\bm{s}\leftarrow\bm{s^{\prime}}$;
+32
+33       end while
+      /* Policy Evaluation */
+34       Periodically evaluate the performance of the current policy;
+35
+36 end for
+return Optimal maintenance action $a$
+Algorithm 4 A Deep Reinforcement Learning Algorithm for Maintenance Action Recommendation
+V Case Study
+In this section, we perform a case study on aircraft engine maintenance using NASA’s C-MAPASS dataset [28] to evaluate our proposed TranDRL framework.
 V-A Dataset
-The NASA’s C-MAPSS dataset [18] is partitioned into four subsets denoted as FD001, FD002, FD003, and FD004, each having a training and test set, reflecting distinct operating and fault modes.
+The NASA’s C-MAPSS dataset [28] contains four subsets denoted as FD001, FD002, FD003, and FD004, each having a training and test set, reflecting distinct operating and fault modes.
 The training data comprises inputs from three operating conditions and 21 sensors’ measurements.
 For training, every engine data point serves as a sample paired with its RUL.
-We applied a piecewise linear degradation model [19] for RUL determinations.
-In testing, the last cycle data point for each engine generally stands as the sample, with actual RUL values included in the dataset.
-V-B Evaluation Metric
-To assess the model’s predictive accuracy, we utilized the Mean Squared Error (MSE), given by
-$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},$ (2)
-which quantifies the average squared differences between predicted $\hat{y}_{i}$ and observed $y_{i}$ values. Additionally, the coefficient of determination, $R^{2}$, defined as
-$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}},$ (3)
-where $\bar{y}$ is the mean of observed values, was employed to measure the proportion of variance explained by the model. A higher $R^{2}$ suggests a better fit to the observed data.
-V-C Remaining Useful Life Prediction
-The Temporal Fusion Transformer (TFT) [20] addresses the limitations of traditional time series forecasting models by offering enhanced interpretability and the capability to deal with multiple prediction horizons.
-It merges the advantages of Recurrent Neural Networks (RNNs) and Transformers [21] into a unified encoder-decoder architecture.
-Encoder: The encoder transforms historical time-series data into a high-dimensional space using a blend of convolutional layers and self-attention mechanisms, thereby learning rich representations of the input.
-Temporal Attention: This specialized module empowers the model to focus on different time-steps variably, facilitating the capture of dynamic temporal patterns and long-range dependencies.
-Decoder: Tailored for multi-horizon prediction, the decoder employs an autoregressive model that recursively uses its previous forecasts as inputs for subsequent steps.
-Static Inputs: Apart from time-dependent features, TFT can also integrate static inputs to provide a comprehensive context for forecasting.
-Interpretable Components: To ensure transparency, TFT integrates mechanisms like gating and linear autoregressive decoders that allow for human-readable prediction interpretation.
-Figure 1: FD001.
-Figure 2: FD002.
-Figure 3: FD003.
-Figure 4: FD004.
-V-D Optimal Maintenance Actions Recommendation
-In the previous stage, we succeed in predicting RUL using four datasets from NASA C-MAPSS.
-But results from the prediction cannot provide enough information for engineers to make decisions on which maintenance actions to take.
-Thus, we may recommend corresponding actions based on the cost and benefits.
-To optimize results, we leverage DRL algorithms, which helps to find the best trade-off.
-V-D1 States and Actions
+In the experiment, we applied a piecewise linear degradation model [29] for RUL determinations.
+The last cycle data point for each engine generally stands as the sample, with actual RUL values included in the dataset.
+Figure 2: An Overview of the Temporal Fusion Transformer.
+V-B Remaining Useful Life Prediction
+The framework given in Algorithm 3 makes it suitable to incorporate the Transformer-based time-series event forecasting networks with other feature extraction and post-processing modules.
+In the case study, we adopt the Temporal Fusion Transformer (TFT) [30] in Algorithm 3 to utilizes the advantages of both recurrent neural networks and Transformers [31] into a unified encoder-decoder architecture.
+Compared to traditional time series forecasting models, e.g., Long Short-Term Memory networks (LSTMs) [32], TFT offers enhanced interpretability and the capability to deal with multiple prediction horizons.
+The structure of TFT is shown in the Figure 2 and the includes following components:
+Encoder. The encoder transforms the raw time-series data into a high-dimensional space using convolutional layers and self-attention mechanisms, thereby learning rich representations of the input.
+Temporal Attention. This specialized module let the model focus on different time-steps, facilitating the capture of dynamic temporal patterns and long-range dependencies.
+Decoder. Tailored for multi-horizon prediction, the decoder employs an autoregressive model that recursively uses its previous forecasts as inputs for forecasting subsequent steps.
+Static Inputs. Apart from time-dependent features, TFT can also integrate static inputs to provide a comprehensive context for forecasting.
+Interpretable Components. To ensure transparency, TFT integrates mechanisms like gating and linear autoregressive decoders that allow for human-readable prediction interpretation.
+Evaluation Results.
+In our comparisons, Figures 3 to 6 demonstrate the TFT algorithm’s superior performance over NBeats [33] and DeepAR [34] in predicting the RUL for a individual machine, as evidenced by consistently higher RUL values.
+This suggests TFT’s enhanced capability in capturing complex, time-dependent patterns in sensor data, highlighting its robustness across various machine behaviors.
+In contrast, Figure 7 shows that while federated learning models generate reasonably good average RUL values across multiple machines, they may not capture the intricate details as effectively as single-machine models.
+However, Figures 8 to 11 show a significant bias when models trained on a single machine are applied to other machines, underscoring the challenge of model generalizability in diverse operational environments.
+This analysis illustrates the trade-off between the granularity and specificity of single-machine models and the broader applicability but potentially lower precision of federated learning approaches, highlighting the need for context-aware modeling strategies in predictive maintenance.
+Figure 3: The RUL prediction result for machine FD001.
+Figure 4: The RUL prediction result for machine FD002.
+Figure 5: The RUL prediction result for machine FD003.
+Figure 6: The RUL prediction result for machine FD004.
+Figure 7: Comparison of RUL prediction results of four machines using the averaged model.
+Figure 8: Comparison of RUL prediction results of four machines using the model trained from machine FD001.
+Figure 9: Comparison of RUL prediction results of four machines using the model trained from machine FD002.
+Figure 10: Comparison of RUL prediction results of four machines using the model trained from machine FD003.
+Figure 11: Comparison of RUL prediction results of four machines using the model trained from machine FD004.
+V-C Optimal Maintenance Actions Recommendation
+In the RUL prediction stage, experiment results show that our proposed framework succeeds in predicting RULs using four datasets from NASA C-MAPSS dataset [28].
+However, a gap still exists between the prediction results and the maintenance action decisions needed by engineers.
+To make action recommendation by weighing the cost and benefits, we leverage DRL algorithms as given in Algorithm 4, which helps to find the best trade-off.
+V-C1 States and Actions
 The state space is generated from PCA features of the RUL resulted from the prediction outcomes of the NASA’s C-MAPSS Aircraft Engine Simulator dataset.
-These PCA features provide a compact yet informative snapshot of the machine’s current health condition.
-The action space contains three primary actions:
+These PCA features provide a compact and important part of the machine’s current health condition.
+The action space contains three actions:
   • No Action ($a_{1}$), appropriate for healthy machine states, minimizes immediate costs but risks penalties in critical conditions;
-  • Partial Maintenance ($a_{2}$), suitable for moderate wear states, involves preventive actions like lubrication to extend machine life at moderate costs;
-  • Complete Overhaul ($a_{3}$), necessary for critical machine states, entails comprehensive repairs and replacements to avert catastrophic failures but incurs high costs and downtime.
+  • Partial Maintenance ($a_{2}$), suitable for moderate unhealthy states, involves preventive actions like lubrication to extend machine life at moderate costs;
+  • Complete Overhaul ($a_{3}$), necessary for critical machine states, entails comprehensive repairs and replacements to prevent catastrophic failures but incurs high costs and downtime.
 The DRL agent aims to navigate this state-action space to minimize long-term operational costs by selecting the optimal maintenance actions based on each state’s representation of machine health.
-V-D2 Probabilities Model
+V-C2 Probabilities Model
 In our study, the transition and action cost matrices are empirically calibrated using the NASA C-MAPSS dataset.
@@
 Moreover, action-related costs, such as maintenance expenses or energy consumption, are grounded in real-world data from this dataset.
-In this MDP framework, the transition matrix has dimensions $(10,3,10)$, representing 10 states, 3 actions, and 10 possible next states, all closely mirroring scenarios that commercial aero-engines might encounter.
-V-D3 Reward Model
-In predictive maintenance applications using DRL, the objective is not only to maximize the RUL of machinery but also to minimize the operational costs and downtime.
+In this MDP framework, the transition matrix has a dimension $(10,3,10)$, representing transitions between $10$ discretized states given $3$ actions, all closely reflecting scenarios that commercial aero-engines might encounter.
+V-C3 Reward Model
+The objective of maintenance actions recommend algorithms is not only to maximize the RUL but also to minimize the operational costs and downtime.
 To fulfill these multi-objective requirements, we propose a reward function that combines these elements in a weighted manner.
@@
 This gain is computed using the state-transition probabilities $p(S_{t+1}|S_{t},a_{t})$ derived from the C-MAPSS dataset.
-A higher RUL gain is desirable as it suggests an extension in the machine’s operational lifespan.
-  • $\text{Cost}(a_{t})$ denotes the operational or maintenance cost incurred by taking action $a_{t}$.
-This cost is modeled based on real-world operation costs and directly affects the overall expenditure.
-  • $\text{DownTime}(a_{t})$ signifies the machine downtime resulting from action $a_{t}$.
-Downtime is critical in operational settings where machine availability significantly impacts productivity or service levels.
-  • $\alpha,\beta,\gamma$ are hyperparameters that weigh the importance of RUL, cost, and downtime, respectively.
-By fine-tuning these hyperparameters, the model can be customized to focus on extending the RUL of machinery, minimizing costs, or reducing downtime, as per the organizational objectives.
-V-D4 Hyperparamters
-To train the algorithm, we use several hyperparameters that were carefully chosen based on empirical experiments. In the following, we present detailed hyperparameters for each DRL algorithm.
+A higher RUL gain is desirable as it suggests an extension in the machine’s normal working time.
+  • $\text{Cost}(a_{t})$ denotes the operational or maintenance cost incurred by taking action $a_{t}$. This cost is modeled based on experiences.
+  • $\text{DownTime}(a_{t})$ signifies the machine downtime resulting from action $a_{t}$. Downtime is critical in operational settings, which impacts productivity or service levels.
+  • $\alpha,\beta,\text{and}~{}\gamma$ are hyperparameters that weigh the importance of RUL, cost, and downtime, respectively.
+By fine-tuning these hyperparameters, the model can be customized to focus on extending the RUL, minimizing costs, or reducing downtime.
+V-C4 Hyperparamters
+For each DRL algorithm embedded in Algorithm 4, we carefully choose its corresponding hyperparameters based on empirical experiments. Details for each DRL algorithm are given as follows.
 Deep Q-network (DQN).
-DQN alleviates the computational burden of tabular Q-learning by employing a deep neural network to approximate the state-action value function.
+DQN reduces the computational burden of tabular Q-learning by employing a deep neural network to approximate the state-action value function [35].
 It also introduces experience replay to improve learning stability and efficiency.
-The learning rate was set to $0.0001$, the discount factor was set to $0.99$, and the replay buffer size was set to $1,000,000$ experiences.
-We used a batch size of $512$ and updated the target network every $50$ steps.
-The exploration vs.
-exploitation trade-off was controlled by an epsilon value that started at $1.0$ and decayed over time to a final value of $0.01$.
-Additionally, we used a neural network with two hidden layers of $256$ units each, and we used the Adam optimizer with a learning rate of $0.0001$.
+The learning rate is set to $10^{-4}$, the discount factor is set to $0.99$, and the replay buffer size is set to $10^{6}$ experiences.
+We use a batch size of $512$ and update the target network every $50$ steps.
+The exploration-exploitation trade-off is controlled by an $\epsilon$ value that started at $1.0$ and decayed over time to a final value of $0.01$.
+Additionally, we use a neural network with two hidden layers of $256$ units each, and we utilize the Adam optimizer with a learning rate of $10^{-4}$.
 Proximal Policy Optimization (PPO).
-PPO ensures more stable and efficient policy updates by employing a clipped objective function, thereby avoiding large, destabilizing updates to the policy.
+PPO ensures more stable and efficient policy updates by employing a clipped objective function, thereby avoiding large, destabilizing updates to the policy [36].
 We set the learning rate for the actor and critic networks to $0.0001$ and the discount factor to $0.99$.
-A clip range of $0.2$ was used to limit the updates to the policy, and we applied a value function coefficient of $0.5$ to balance the importance of the policy and the value function.
-Additionally, we set the entropy coefficient to $0.01$ to encourage exploration, and we used a batch size of $64$.
-We optimize the policy every 10 epochs and used the Adam optimizer with an epsilon value of $0.00001$.
-Finally, we set the GAE parameter to $0.95$ to estimate the advantage function.
+A clip range of $0.2$ is used to limit the updates to the policy, and we apply a value function coefficient of $0.5$ to balance the importance of the policy and the value function.
+Additionally, we set the entropy coefficient to $0.01$ to encourage exploration and a batch size as $64$.
+Then, we optimize the policy every 10 epochs and use the Adam optimizer with a learning rate of $0.0001$.
+Finally, GAE parameter is set to $0.95$ to estimate the advantage function.
 Soft Actor-Critic (SAC).
-SAC is particularly effective for problems with high-dimensional action spaces.
+SAC is particularly effective for problems with high-dimensional action spaces [37].
 SAC aims to maximize an objective function that balances both the expected return and the entropy of the policy, making it more robust and efficient.
-The learning rate for the actor and critic networks was set to $0.0003$, and we used a discount factor of $0.99$.
-We applied a temperature parameter alpha of $0.2$ to balance the entropy term in the objective function.
-The target entropy was set to $-2.0$, and we used a replay buffer of $1,000,000$ experiences with a batch size of $256$.
-We also applied a soft update of the target networks with a rate of $0.005$, and we initialized the log standard deviation of the policy with a value of $-2$.
-We trained the algorithm for $100,000$ time steps and used the Adam optimizer with an epsilon value of $0.0001$.
-V-E Results and Discussion
-We compare the performance of DQN [22], PPO [23], and SAC [24] on our designed task in the Figure 5, Figure 6, and Figure 7.
+The learning rate for the actor and critic networks is set to $0.0003$, and we use a discount factor of $0.99$.
+Then, we apply a temperature parameter alpha of $0.2$ to balance the entropy term in the objective function.
+The target entropy is set to $-2.0$, and we use a replay buffer of $10^{6}$ experiences with a batch size of $256$.
+We also apply a soft update of the target networks with a rate of $0.005$, and we initialize the log standard deviation of the policy with a value of $-2$.
+Finally, we train the algorithm for $10^{5}$ time steps and use the Adam optimizer with a learning rate of $0.0001$.
+V-D Results and Discussion
+Figure 12: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
+Figure 13: The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
+We compare the performance of DQN, PPO, and SAC on our designed task in Figures 1213.
 The evaluation metric is the total rewards obtained by the agent during a fixed number of episodes.
@@
 On the other hand, if the action space is continuous or the observations are high-dimensional, PPO and SAC may be more suitable and converge faster.
-Results that are shown in Figure 5 and Figure 7 have confirmed this theory.
 Overall, these experimental results demonstrate the importance of selecting the appropriate DRL algorithm and hyperparameters for a specific task to achieve the optimal performance.
-Figure 5: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
-Figure 6: The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
-Figure 7: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms in Cartpole environment.
+Figure 14: Differences between RL and RLHF.
+V-E Reinforcement Learning with Human Feedback
+Incorporating human expertise into the reinforcement learning process helps to resolve some of the challenges encountered with traditional RL methods.
+Especially, in scenarios where real-world rewards are hard to quantify or can be misleading, human feedback provides a way to guide the learning process more effectively.
+The comparison between RL and RLHF is presented in Fig.
+14.
+As machines and their associated systems become increasingly complex, the conditions that need to be considered maybe take a long time for Algorithm 5.
+Human experts, with their working experiences and domain knowledge to find complex patterns, can provide valuable insights that may not be readily available through raw data.
+By integrating human feedback into the RL process, the framework can:
+  • Augment the learning process with domain-specific knowledge.
+  • Address the cold start problem where initial exploration might be risky or expensive.
+  • Improve learning efficiency by reducing the exploration of sub-optimal actions.
+Thus, we further incorporate the following information into our proposed framework in Algorithm 5:
+  1. 1.
+    Human Opinions: Collect human predictions of the desired behavior. Experienced engineers can provide opinions on maintenance decisions based on various machine states from the C-MAPSS dataset.
+  2. 2.
+    Feedback on Agent’s Actions: After the agent makes a decision, a human can provide feedback on the chosen action, guiding the agent on the appropriateness of its actions.
+Then, we introduce a term in the reward function that captures human feedback
+$R_{\text{human}}(a_{t})=\begin{cases}r_{\text{positive}},&\text{if human feedback is positive},\\ r_{\text{negative}},&\text{if human feedback is negative},\\ 0,&\text{otherwise},\end{cases}$ (7)
+based on which, the modified reward function for DRL is
+$R(S_{t},a_{t},S_{t+1})$ $=\alpha\times\text{RUL}_{\text{gain}}(S_{t},a_{t})-\beta\times\text{Cost}(a_{t})-$
+$\gamma\times\text{DownTime}(a_{t})+\delta\times R_{\text{human}}(a_{t}),$
+where $\delta$ is a hyperparameter weighing the importance of human feedback.
+Input: Observation $\bm{s}$, Exploration rate $\epsilon$, Learning rate $\alpha$, Discount factor $\gamma$, Human feedback 00.
+Output: Optimal action $a$ for maintenance
+1 Initialize the neural network weights;
+2 Initialize a mechanism to incorporate human feedback (e.g., reward shaping, action recommendation);
+3 for each episode do
+       /* Incorporate Human Demonstrations */
+4       If available, utilize human demonstrations to set initial policy;
+5       Initialize the starting state $\bm{s}$;
+6       while $s$ is not a terminal state do
+             /* Action Selection with Human Feedback */
+7             Select action $a$ using an $\epsilon$-greedy policy or based on human recommendations;
+8             Execute action $a$, observe reward $r$ and successor state $\bm{s^{\prime}}$;
+             /* Adjust Reward Based on Human Feedback */
+9             Incorporate human feedback into the reward:
+$R_{\text{human}}(a_{t})=\begin{cases}r_{\text{positive}},&\text{if HF is positive},\\ r_{\text{negative}},&\text{if HF is negative},\\ 0,&\text{otherwise}.\end{cases}$
+Adjust reward $r$ based on $R_{\text{human}}(a_{t})$;
+             /* Network Update */
+10             Calculate TD error $\delta=r+\gamma\max_{a^{\prime}}Q(\bm{s^{\prime}},a^{\prime})-Q(\bm{s},a)$;
+11             Update neural network weights using $\delta$ and learning rate $\alpha$;
+             /* Adjustment of Exploration and State */
+12             Adjust the exploration rate $\epsilon$;
+13             Update current state $\bm{s}\leftarrow\bm{s^{\prime}}$;
+14
+15       end while
+      /* Incorporate Feedback into Learning Process */
+16       Periodically adjust the learning process based on accumulated human feedback;
+17
+18 end for
+return Optimal maintenance action $a$.
+Algorithm 5 The RLHF Algorithm for Maintenance Actions Selection
+V-F Results for RLHF
+Figure 15: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
+Figure 16: The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
+Figure 15 and Figure 16 reflect the changes after adding on the human feedback in the reward function.
+The results show that the speed of convergence does not improve too much, but the process becomes more stable and variances minimized.
+This is affected by the weight of human feedback in the reward function as well as the complexity of the task.
+Since our task converges very fast even without human feedback, there is not much room for improvement.
+In the future, we will conduct experiments on more complex real-world scenarios to confirm the effectiveness of the RLHF algorithm.
 VI Conclusions
-In this paper, we have presented a comprehensive prescriptive maintenance framework that uniquely integrates real-time RUL estimations with reinforcement learning algorithms for dynamic action recommendations.
-The framework starts by employing the transformer algorithm to estimate the RUL distribution of engines, offering a robust, adaptive approach to understanding machine health in real-time.
-As machinery operates, our model continues to dynamically update these RUL estimates, providing an ever-evolving understanding of the equipment’s condition.
-We rigorously compared the performance of several state-of-the-art time-series forecasting models, including TFT, NBeats, and DeepAR, to validate the accuracy of our RUL estimates.
+In this paper, we present a comprehensive prescriptive maintenance framework that integrates real-time RUL estimations with DRL algorithms for dynamic action recommendations.
+The framework starts by employing the Transformer network to estimate the RUL distribution of engines, offering a robust and adaptive data-driven approach to understanding machine conditions in real-time.
+As machine operates, our model continues to dynamically updating RUL estimates, providing an ever-evolving status of the equipment’s condition.
+We rigorously compare the performance of several state-of-the-art time-series forecasting models, including TFT, NBeats, and DeepAR, to validate the accuracy of our RUL estimates.
 Our findings indicate that the Transformer-based algorithm delivers superior predictive performance, laying a solid foundation for the subsequent decision-making process.
-For recommending maintenance actions, we used SAC, DQN, and PPO.
-Each of these algorithms has its strengths and weaknesses, but they all were adapted to work seamlessly with the dynamically updated RUL estimates.
-Our experimental results showed that SAC excelled in scenarios where the cost function was highly non-linear, while PPO and DQN were more suited for environments with more predictable dynamics.
-This framework opens up new avenues for research, particularly in the fine-tuning of hyperparameters that govern the trade-offs between extending the machinery’s useful life and minimizing operational costs.
-Future work could involve integrating more complex cost models, incorporating uncertainty estimates, and testing the framework in a multi-agent setting or across different types of machinery.
-In conclusion, our work signifies a step forward in the quest for more intelligent, adaptive, and cost-effective prescriptive maintenance systems.
-By leveraging cutting-edge algorithms for both RUL estimation and action recommendation, we offer a holistic solution that dynamically adapts to the operational realities, thereby driving both economic and operational efficiencies.
+For DRL-based maintenance action recommendation, we demonstrate the efficiency of SAC, DQN, and PPO.
+Each of these algorithms has its strengths and weaknesses, but they all are adapted to working seamlessly with the dynamically updated RUL estimates.
+Future work may involve integrating more complex cost models, incorporating uncertainty estimates, and testing the framework in a multi-agent setting or across different types of machines.
+In conclusion, our work signifies a step forward in the request for more intelligent, adaptive, and cost-effective prescriptive maintenance systems.
+By leveraging cutting-edge algorithms for both RUL estimation and action recommendation, we offer a solution that dynamically adapts to the operational realities, thereby enhancing both economic and operational efficiencies.

### v3

@@
 Abstract
-Industrial systems demand reliable predictive maintenance strategies to enhance operational efficiency and reduce downtime.
-This paper introduces an integrated framework that leverages the capabilities of the Transformer model-based neural networks and deep reinforcement learning (DRL) algorithms to optimize system maintenance actions.
-Our approach employs the Transformer model to effectively capture complex temporal patterns in sensor data, thereby accurately predicting the remaining useful life (RUL) of an equipment.
+Industrial systems require reliable predictive maintenance strategies to enhance operational efficiency and reduce downtime.
+Existing studies rely on heuristic models which may struggle to capture complex temporal dependencies.
+This paper introduces an integrated framework that leverages the capabilities of the Transformer and Deep Reinforcement Learning (DRL) algorithms to optimize system maintenance actions.
+Our approach employs the Transformer model to effectively capture complex temporal patterns in IoT sensor data, thus accurately predicting the Remaining Useful Life (RUL) of equipment.
 Additionally, the DRL component of our framework provides cost-effective and timely maintenance recommendations.
-We validate the efficacy of our framework on the NASA C-MPASS dataset, where it demonstrates significant advancements in both RUL prediction accuracy and the optimization of maintenance actions, compared to the other prevalent machine learning-based methods.
-Our proposed approach provides an innovative data-driven framework for industry machine systems, accurately forecasting equipment lifespans and optimizing maintenance schedules, thereby reducing downtime and cutting costs.
+Numerous experiments conducted on the NASA C-MPASS dataset demonstrate that our approach has a performance similar to the ground-truth results and could be obviously better than the baseline methods in terms of RUL prediction accuracy as the time cycle increases.
+Additionally, experimental results demonstrate the effectiveness of optimizing maintenance actions.
 Index Terms:
@@
 I Introduction
-The fourth industrial revolution, colloquially known as Industry 4.0, has ushered in unprecedented levels of automation and data exchange in manufacturing technologies [1].
-With these advancements, however, comes the increased requests for reliable, efficient, and cost-effective maintenance systems.
-Prescriptive maintenance has thus emerged as a vital area of research and industrial application, offering the promise of proactive fault detection and optimized action planning [2].
+Industry 4.0 has introduced unprecedented levels of automation and data exchange in manufacturing technologies [1].
+However, with these advancements comes increased requirement for reliable, efficient, and cost-effective maintenance systems.
+Prescriptive maintenance has thus emerged as a crucial topic in both research and industrial sectors, offering the promise of proactive fault detection and optimized operation strategy planning [2].
 It differs from reactive maintenance, which only kicks in post-failure, causing often unplanned and expensive downtimes.
@@
 However, current predictive maintenance systems largely rely on heuristic models for action planning or conventional machine learning algorithms for sensor data processing.
-While effective to some extent, existing approaches often fail to capture complex temporal dependencies in sensor data due to issues like vanishing or exploding gradients, limiting their ability to accurately predict Remaining Useful Life (RUL).
-Apart from that, there is a trend in recent works to leverage Reinforcement Learning (RL) to make plans in maintenance tasks [4].
-RL excels at making optimal decisions in complex environments, offering a powerful tool for action planning in maintenance tasks.
-As a result, the RL-based decisions offer optimal maintenance action schedules that balance the trade-off between equipment reliability and maintenance costs.
+Although effective to some extent, existing approaches often fail to capture complex temporal dependencies in sensor data due to issues such as vanishing or exploding gradients, limiting their ability to accurately predict RUL.
+To address these issues, a number of new frameworks (e.g., attention in recurrent networks [4]) are developed to capture the correlation in temporal sequence data.
+Apart from that, there is a trend in recent studies to leverage DRL to make plans in maintenance tasks [5].
+DRL excels at making optimal decisions in complex environments, offering a powerful tool for action planning in maintenance tasks.
+As a result, the DRL-based decisions offer optimal maintenance action schedules that balance the trade-off between equipment reliability and maintenance costs.
 However, these methods consistently encounter the cold start problem [6], coupled with low learning efficiency.
-In some special cases such as aeroengine, these methods also suffer from a lack of reliability and explainability.
+In some special cases, such as aero-engines, these methods also suffer from a lack of reliability and explainability.
 This often leads to their decisions being perceived as untrustworthy by users.
-Unlike existing solutions that address either predictive or prescriptive maintenance, we propose a framework named TranDRL, which integrates both aspects seamlessly.
-TranDRL leverages advanced machine learning techniques to not only accurately predict RUL but also to convert these predictions into actionable maintenance recommendations.
-By integrating these two critical components, our framework provides a comprehensive solution for industrial maintenance challenges by Transformer neural networks and Reinforcement Learning with Human Feedback (RLHF) algorithms.
-Transformers excel at modeling long-term dependencies in sequential data by the attention mechanism and are highly adaptable, making them ideal for handling complex industrial sensor data.
-By leveraging the expertise and insights of human operators or maintenance professionals, the algorithm can efficiently guide its actions, accelerating the convergence of the DRL algorithm towards optimal maintenance strategies.
-This human-guided approach ensures that the algorithm remains on the right track, continually improving its performance and ultimately leading to more efficient and cost-effective industrial maintenance practices.
-Contributions. The contributions of this paper can be summarized as follows:
-  • To improve the accuracy of the RUL prediction module, we adopt Transformer to address the difficulty in reflecting long-term dependencies of sequential sensor data in industrial scenarios.
-  • We propose a DRL-based recommendation system for determining maintenance actions, reducing both labor involvement and operational costs.
-Furthermore, we utilize RLHF to confirm the correct maintenance actions and improve learning efficiency.
-  • A Transformer-DRL framework is proposed to optimize the prescriptive maintenance process, which can be extended to multiple manufacturing scenarios as well.
-  • Extensive simulation experiment results, conducted on different datasets, demonstrate that our method outperforms baseline methods in terms of both predicting RUL and optimizing maintenance actions.
+Unlike existing solutions, we propose a framework named TranDRL, which integrates both aspects seamlessly.
+TranDRL leverages both the Transformer framework and DRL to not only accurately predict the RUL but also convert these predictions into maintenance action recommendations.
+Additionally, we adopt the RL with Human Feedback (RLHF) algorithm, which incorporates the human expert’s opinion on maintenance actions into the DRL algorithm to facilitate the algorithm in making decisions.
+By integrating these two critical neural networks and human expert’s experience, our framework provides a comprehensive solution for industrial maintenance challenges.
+Transformers excel at modeling long-term dependencies in sequential data through their attention mechanism and exhibit a high degree of adaptability, making them ideal for handling complex industrial sensor data.
+Utilizing the expertise and insights of human maintenance professionals, the algorithm can efficiently guide the production operational actions, expediting the convergence of the DRL algorithm toward optimal maintenance strategies.
+This human-guided approach ensures that the algorithm remains on the right track, continually improving its performance and leading to more efficient and cost-effective industrial maintenance practices.
+Contributions. The main contributions of this paper can be summarized as follows:
+  • To improve the accuracy of the RUL prediction module, we adopt a Transformer-based neural network to address the difficulty in reflecting long-term dependencies of sequential sensor data.
+  • We propose a DRL-based recommendation approach for determining maintenance actions, reducing both labor involvement and operational costs.
+Furthermore, we utilize the RLHF algorithm to confirm the correct maintenance actions and improve the learning efficiency of the DRL framework.
+  • The transformer-enabled DRL algorithm is tailored into a federated learning framework, which can optimize the prescriptive maintenance process in multi-machine scenarios.
+  • Extensive simulation experiment results demonstrate that our method outperforms a series of baseline methods.
 Organizations.
@@
 Section II reviews related work in the field of prescriptive maintenance utilizing machine learning algorithms.
-Section III formulate the problem we are trying to solve.
+Section III formulates the problem we are trying to solve.
 Then, Section IV describes our proposed framework for Transformer-driven prediction maintenance in detail.
@@
 Finally, Section VI concludes the paper and outlines future directions for research.
-Notations used across the paper is provided in Table I.
-TABLE I: Notations.
-Symbol Description
-$\mathbf{X}$
-Time series of sensor readings and operational parameters.
-$R(t)$
-Remaining Useful Life at time $t$: $R(t)=T-t$.
-$T$
-Total expected lifespan of the component.
-$t$
-Current operational time of the machine.
-$M$
-Model estimating $R(t)$ from $\mathbf{X}$.
-$s$
-State encapsulating health and RUL metrics.
-$\mathcal{A}$
-Set of maintenance actions.
-$a$
-Action chosen to maximize future reward.
-$a^{*}$
-Optimal action maximizing cumulative benefit.
-$Q(s,a)$
-Expected reward of action $a$ in state $s$.
-$\epsilon$
-Exploration rate in DRL.
-$\alpha$
-Learning rate in DRL.
-$\gamma$
-Discount factor for future rewards.
-$\bm{y}$
-Predicted RUL output by Transformer model.
-$\bm{s}$
-Observation state in DRL from RUL predictions.
-$d_{\text{model}}$
-Input/output dimension in Transformer.
-$d_{k},d_{v}$
-Dimensions of keys/queries and values in attention.
-$\mathbf{Q},\mathbf{K},\mathbf{V}$
-Query, key, and value matrices in attention.
-$FFN(\cdot)$
-Feed-forward network in Transformer.
-$\bm{\Lambda}$
-Attention weight matrix.
-$\alpha_{ij}$
-Attention weight from $j$-th input to $i$-th output.
-$\bm{h}$
-Hidden states in encoder.
-O
-Output states from decoder.
-$\sigma(\cdot)$
-Activation function in prediction module.
-$K$
-Number of machines in federated learning setup.
-$f_{\text{ensemble}}$
-Function for RUL prediction in multi-machine case.
-$\mathbf{W},b$
-Weight matrix and bias in prediction module.
-PCA
-Used for selecting important parts of RUL.
-Figure 1: An Overview of TranDRL Framework.
+Notations used across the paper are provided in Table I.
+TABLE I: Notations and abbreviations.
+Symbol Descriptions
+$\mathbf{X}$ Time series of sensor readings and operational parameters.
+$R(t)$ Remaining Useful Life at time $t$: $R(t)=T-t$.
+$T$ Total expected lifespan of the component/machine.
+$t$ Current operational time of the machine in concern.
+$M$ Model estimating $R(t)$ from $\mathbf{X}$.
+$s$ State encapsulating the health and RUL metrics.
+$\mathcal{A}$ Set of maintenance actions.
+$a$ Action chosen to maximize the future reward.
+$a^{*}$ Optimal action maximizing the cumulative benefit.
+$Q(s,a)$ Expected reward of action $a$ in state $s$.
+$\epsilon$ Exploration rate in DRL.
+$\alpha$ Learning rate in DRL.
+$\gamma$ Discount factor for future rewards.
+$\bm{y}$ Predicted RUL output by the neural network model.
+$\bm{s}$ Observation state in DRL from RUL predictions.
+$d_{\text{model}}$ Input/output dimension of the Transformer.
+$d_{k},d_{v}$ Dimensions of keys/queries and values in attention computation.
+$\mathbf{Q},\mathbf{K},\mathbf{V}$ Query, key, and value matrices in attention computation.
+$FFN(\cdot)$ Feed-forward network in the Transformer module.
+$\bm{\Lambda}$ Attention weight matrix.
+$\alpha_{ij}$ Attention weight from the $j$-th input to the $i$-th output.
+$\bm{h}$ Hidden states in the encoder.
+O Output states from the decoder.
+$\sigma(\cdot)$ Activation function in the prediction module.
+$K$ Number of machines in federated learning setups.
+$f_{\text{ensemble}}$ Function for RUL prediction in the multi-machine case.
+$\mathbf{W},b$ Weight matrix and bias in the prediction module.
+PCA Principal component analysis.
+DQN Deep Q-network.
+PPO Proximal Policy Optimization.
+SAC Soft Actor-Critic.
 II Related Work
-[…]
-Corrective maintenance is “fault-based maintenance” and is an unplanned mode of maintenance carried out only when equipment malfunctions are detected.
-Preventive maintenance, or scheduled inspection and maintenance, generally involves halting the production line according to a predetermined schedule, with an emphasis on repairs before a failure occurs.
-These two types of maintenance methods are traditionally the most commonly adopted in the industry.
-However, they often lead to production losses due to their neglect of the historical operational status of the production system and components, resulting in either too frequent or too infrequent maintenance.
-Predictive maintenance determines the timing and methods of maintenance based on the physical model of the production system and historical monitoring data, and is therefore also referred to as condition-based maintenance.
-Prescriptive maintenance is generally considered an extension of predictive maintenance, optimizing maintenance strategies based on fault diagnosis results.
-It emphasizes identifying the inherent reasons for the occurrence of faults and subsequently directing the generation of operational and maintenance strategies [7].
-Compared to traditional maintenance methods, prescriptive maintenance leverages historical system status to comprehensively forecast future system operations and optimize the resources required for maintenance, aiming to prevent waste from excessive maintenance or production losses due to malfunctions and downtime.
+Prescriptive maintenance uses historical operational data and (possibly implicit) system models to optimize maintenance timing and strategies.
+It extends the prediction maintenance approaches by optimizing strategies based on fault diagnosis, focusing on identifying root causes and selecting proper maintenance actions [7].
+Compared to traditional methods, it aims to prevent waste and production losses by leveraging historical system status to forecast future machine status and optimize maintenance resources.
 II-A Deep Learning-based Remaining Useful Life Prediction
 In production maintenance tasks, accurately predicting the RUL of machinery has traditionally required intricate models that capture the nuances of production system operations.
-While many revert to statistical signal processing and machine learning to bypass comprehensive system modeling, these methods often mandate profound domain-specific knowledge, especially for tasks like signal denoising and feature extraction [8, 9].
+While many revert to statistical signal processing and machine learning to bypass comprehensive system modeling, these methods often require deep domain-specific knowledge, especially for tasks such as signal denoising and feature extraction [8, 9].
 Notably, they also entail labor-intensive manual feature engineering.
@@
 Deep learning, boosted by computational advances, emerges as a remedy.
-Pioneering algorithms, such as Convolutional Neural Networks (CNN) and Deep Autoencoders, have gained traction for their aptitude in fault diagnosis and predictive maintenance [10, 11].
-Their intrinsic capability to distill hierarchical features from raw data eliminates manual feature engineering, fostering accurate RUL predictions.
+Pioneering algorithms, such as Convolutional Neural Networks (CNN) and Deep Autoencoders, have gained popularity owning to their proficiency in fault diagnosis and predictive maintenance [10, 11].
+Their intrinsic capability to distill hierarchical features from raw data eliminates the need of manual feature engineering, thus fostering more accurate RUL predictions.
 Amongst the deep learning architectures, Transformers, originally conceived for language processing tasks [12], have recently demonstrated efficacy in time-series data processing.
 For instance, Zhou et al.
-[13] integrated a time-embedding layer, augmenting the Transformer’s time-series prediction.
+[13] integrated a time-embedding layer, augmenting the Transformer’s time series prediction.
 Yet, a full exploration of Transformers in the realm of prescriptive maintenance remains on the horizon.
-II-B Deep Reinforcement Learning enabled Predictive Maintenance
-DRL’s concept into predictive maintenance has primarily focused on refining maintenance schedules [14, 2, 4].
+II-B Deep Reinforcement Learning-enabled Predictive Maintenance
+DRL’s concept of predictive maintenance has primarily focused on refining maintenance schedules [14, 2, 5].
 While these applications exhibit DRL’s capability in mitigating machinery breakdowns, they predominantly center on forecasting malfunctions, focusing on actionable preventative measures.
-Gordon et al.
-[15] explore into employing DRL for prescriptive maintenance, but the integration of prediction and actionable recommendations remained elusive.
-In existing maintenance studies, we have explored a lot at preventive and predictive methods, as mentioned in [16] and [17].
-But now, there is growing interest in prescriptive maintenance, which goes beyond just predicting problems to suggesting the best actions to take.
-Even though the work in [16] and [17] has started to light the way, most efforts are either focused on predicting what will happen or on prescribing what to do.
-Even with improvements in using Transformer technology for predictions and DRL for making recommendations, there is no plan to integrate these two ideas together.
-Our study aims to fill this gap by introducing a way to use both the forecasting power of Transformers and the decision-making abilities of DRL to not only predict equipment failures but also suggest the best actions to prevent them.
-II-C Deep Reinforcement Learning with Human Feedback
-Inspired by preference theory [18], Reinforcement Learning from Human Feedback (RLHF) improves the performance of DRL by leveraging feedbacks from human, and releases the difficulty of designing reward function by incorporating human guidance and expertise [19].
-To tackle the problem of informative limitation and expensive time consumption, Ibarz et al.
-[20] combine expert demonstrations and trajectory preferences together and attempt to let human interact with agent directly.
-To enhance sample efficiency, other follow-up works leverage semi-supervised learning and data augmentation techniques into RLHF [21, 22].
-Most previous works are under the non-Markov assumption which poses a threat to modeling reward function, Early et al.
-[23] release this assumption and formulate the scenario into a multiple instance learning process.
-With the prosperity of large language model recently, RLHF also shows its power in fine-tuning step before the model development in this specific domain [24, 25].
-Furthermore, in safety-critical applications like aeroengine maintenance, deviating from standard procedures can have grave consequences.
-Such non-standard maintenance practices might result in serious safety incidents.
-Combining human feedback can also help in setting safe boundaries and ensuring that the agent’s behavior remains within acceptable limits.
-To make the policy more reliable, the incorporation of human feedback can make the agent’s behavior more predictable and better aligned with human expectations, enhancing the explainability of its actions and building trust among users.
+Recent work such as [15] explored the possibility of employing DRL for prescriptive maintenance, but the integration of prediction and actionable recommendations remained elusive.
+In existing maintenance studies, researchers have explored a number of preventive and predictive methods, as mentioned in [16] and [17].
+However, there is a growing interest in prescriptive maintenance, which extends beyond merely predicting problems, instead offering suggestions for the best actions to take.
+Studies such as [16] and [17] have started to light the way, with most efforts focused on either predicting what will happen or prescribing what to do.
+However, even with improvements in using the Transformer technology for predictions and DRL for making recommendations, there is yet no plan to integrate these two ideas.
+Our study aims to fill this gap by introducing a way to utilize both the forecasting power of Transformers and the decision-making capabilities of DRL to not only predict equipment failures but also suggest the best actions to prevent them.
+Furthermore, inspired by preference theory [18], RLHF improves DRL performance by leveraging human feedback, reducing the difficulty in designing reward functions [19].
+To address limitations of information and expensive data collection, expert demonstration and trajectory preferences are combined in [20], enabling direct human-agent interaction.
+Subsequent studies enhanced sample efficiency through semi-supervised learning and data augmentation [21, 22].
+In [23] the non-Markov assumption is relaxed, formulating the scenario as a multiple-instance learning process.
+In safety-critical applications such as aero-engine maintenance, human feedback can set safe boundaries, therfore ensuring that the behavior of an agent remains within acceptable limits, making it more predictable and aligned with human expectations, as well as enhancing explainability and trust.
 III Problem Formulation
@@
 However, with increased automation and reliance on these machines, any unplanned downtime can be a critical blow to productivity and financial stability.
-The cascade effect is not just limited to production loss; it also affects aspects like increased costs for emergency repairs, contractual penalties due to delivery delays, and even reputation damage [26].
-Therefore, the primary challenge is to recommend proper maintenance actions effectively based on accurate prediction of failures of the machine.
-III-A Predicting Remaining Useful Life (RUL)
+The cascade effect is not just limited to production loss; it also affects aspects such as increased costs for emergency repairs, contractual penalties due to delivery delays, and even reputation damage [24].
+Therefore, the primary challenge is to recommend maintenance actions based on accurate machine failure prediction effectively.
+III-A Remaining Useful Life Prediction
 Predicting RUL is essential for predictive maintenance, which provides an estimation of the duration that a machine or its components will continue to function before encountering a failure.
-An accurate RUL prediction enable the industries to strategically schedule their maintenance activities.
+An accurate RUL prediction enables industries to strategically schedule their maintenance activities.
 The RUL offers a dynamic and predictive window into the operational health of a machine.
-The operational data of a machinery component is typically represented as a time series $\mathbf{X}=(\bm{x_{1}},\bm{x_{2}},\ldots,\bm{x_{n}})$, where each $\bm{x}_{i}$ can represent multiple sensor readings, operational parameters, or environmental conditions.
+The operational data of a machinery component is typically represented as a time series $\mathbf{X}=[\bm{x_{1}},\bm{x_{2}},\ldots,\bm{x_{n}}]$, where each $\bm{x}_{i}$ can represent multiple sensor readings, operational parameters, or environmental conditions at a sampling instance $i$.
 The RUL at any given time $t$, denoted by $R(t)$, can be mathematically represented as: $R(t)=T-t$, where
   • $T$ is the total expected operational lifespan of the machine or component, which can vary based on its usage, environmental factors, and maintenance history.
-  • $t$ represents the current age or operational time of the machine.
-The inherent complexity lies in the fact that $T$ is not a constant but a variable that might change based on multiple factors, for example, machine conditions.
-The primary challenge is to develop a predictive model $M$ that can harness the rich time-series data $\mathbf{X}$ to provide an accurate approximation of $R(t)$: $M(\mathbf{X})\approx R(t)$.
-III-B Optimal Maintenance Action Recommendations with DRL
-While predicting the RUL offers industries a foresight into potential downtimes, it is the actionable insights derived from these predictions that truly empower industries to make informed decisions.
+  • $t$ represents the machine’s current age or operational time.
+The inherent complexity is that $T$ is not a constant but a variable that might change based on multiple factors such as machine conditions.
+The primary challenge is to develop a predictive model $M$ that can use the time-series data $\mathbf{X}$ to provide an approximation of $R(t)$: $M(\mathbf{X})\approx R(t)$.
+III-B Optimal Maintenance Action Recommendation
+While predicting the RUL can help industries predict potential downtimes, it is the actionable insights derived from these predictions that empower industries to make informed decisions.
 Upon gaining prediction results of potential machine downtimes, the challenge shifts to determining the most optimal maintenance action.
-Leveraging the capabilities of DRL, our framework not only predicts the machinery’s future status but also suggests the best course of action to prolong its life and ensure optimal performance.
+Leveraging the capabilities of DRL, our framework not only predicts the machinery’s future status but also suggests the best course of action to extend its working time and ensure optimal performance.
 Consider a machinery component’s state $s$, which encapsulates its health, RUL, and other relevant metrics.
 From the state, we want to choose an action $a$ from a set of potential actions set $\mathcal{A}$ to maximize the expected future reward.
-The essence of this problem is rooted in the principle of maximizing long-term rewards.
-We denote a value function $Q(s,a)$, which characterizes the expected cumulative benefit or reward of adopting action $a$ in state $s$.
-The core objective is to navigate the action space and select: $a^{*}=\arg\max_{a\in\mathcal{A}}Q(s,a)$, where
-  • $Q(s,a)$ encompasses immediate benefits and projected future rewards, taking into consideration the dynamic nature of machinery health and operation.
-  • The cumulative reward should integrate both current and weighted future’s benefits strategically.
+Therefore, the challenge of this problem is to maximize delayed reinforcements.
+We both present and denote a value function $Q(s,a)$ that expresses the sum of the utility or profit from choosing $a$ in $s$.
+The objective is to navigate the action space and select: $a^{*}=\arg\max_{a\in\mathcal{A}}Q(s,a)$, where
+  • $Q(s,a)$ reflects immediate benefits and projected future rewards, considering the dynamics of machinery health and operation.
+  • The cumulative reward should integrate the current and the weighted future benefits.
+Figure 1: An Overview of TranDRL Framework.
 IV Proposed Framework
-To solve the problem above, we propose a framework consisting of two major components: a transformer-based learning model to predict the RUL and a decision module to recommend maintenance actions as illustrated in Figure 1 and Algorithm 1.
-Specifically, we consider two cases for the RUL prediction, including a single machine case and a multi-machine case.
+To solve the problem above, we propose a framework consisting of two major components: a transformer-based learning model to predict the RUL and a decision module to recommend maintenance actions, as illustrated in Algorithm 1 and Figure 1.
+Specifically, the sensor data is collected and utilized to predict the RUL of machines, after which DRL algorithms are employed to determine the optimal action for repairing machines.
 Input: Sequence of sensor readings $\mathbf{X}=(\bm{x_{1}},\bm{x_{2}},\ldots,\bm{x_{n}})$, exploration rate $\epsilon$, learning rate $\alpha$, discount factor $\gamma$
 Output: Recommended maintenance action $a$
+1
 /* Phase 1: RUL prediction using Transformer framework */
@@
 4 Generate RUL prediction $\bm{y}=(y_{1},y_{2},\ldots,y_{m})$;
+5
 /* Phase 2: Maintenance action recommendation using DRL */
@@
 return $a$
-Algorithm 1 A Transformer-enabled DRL framework for maintenance selection.
+Algorithm 1 TranDRL:A Transformer-enabled DRL framework for optimal maintenance action recommendation.
 IV-A Case 1: Remaining Useful Life Prediction Module for a Single Machine
-Transformers, rooted in natural language processing [27], have become valuable tools in time-series data analysis.
-This is largely due to their proficiency in understanding complex temporal sequences.
-Our model leverages the Transformer’s capabilities to forecast the RUL from historical time-series sensor data.
-The core of this approach is the self-attention mechanism, which establishes causal relations between different time steps based on query-and-key similarities, adapting to the complexity of the data, for example, self-attention can handle complex relationships in machinery sensor data when issues depend on small changes over time in predictive maintenance.
+Transformers, originating from natural language processing [25], have become valuable tools in time-series data analysis, due to their proficiency in understanding complex temporal sequences.
+Our proposed Algorithm 2 leverages a transformer to forecast the RUL from historical time-series sensor data.
+The core of this approach is the self-attention mechanism, which establishes causal relations between different time steps based on query-and-key similarities, adapting to the complexity of the data, for example, self-attention can handle complex relationships in sensor data while issues depend on small changes over time in predictive maintenance.
 IV-A1 Encoder Module
 For a given time-series $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]$, where $n$ is the sequence length and $\bm{x}_{i}\in\mathbb{R}^{d_{\text{model}}}$, the Encoder module converts $\mathbf{X}$ into hidden states $\bm{h}=[h_{1},h_{2},\cdots,h_{n}]$, with each $h_{i}\in\mathbb{R}^{d_{\text{model}}}$.
-Each Transformer block in this module comprises:
+This module comprises an input embedding layer and a number of $L$ identical self-attention blocks as follows:
+  • Input Embedding.
+The input data in $\mathbf{X}$ are processed with a learned embedding layer and converted to vectors of dimension $d_{\text{model}}$.
+The embedding results are then added with a learned positional encoding of the same dimension to reflect the temporal difference between the input samples.
   • Multi-head Self-attention Mechanism.
-Given matrices for query $\mathbf{Q}\in\mathbb{R}^{n\times d_{k}}$, key $\mathbf{K}\in\mathbb{R}^{n\times d_{k}}$, and value $\mathbf{V}\in\mathbb{R}^{n\times d_{v}}$, the attention mechanism is:
-    $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V},$
-    with $d_{k}$ being the dimension of queries and keys, and $d_{v}$ the dimension of values.
-  • Feed-Forward Neural Network. To capture non-linear patterns, this is represented as:
+Given the matrices for query $\mathbf{Q}\in\mathbb{R}^{n\times d_{k}}$, key $\mathbf{K}\in\mathbb{R}^{n\times d_{k}}$, and value $\mathbf{V}\in\mathbb{R}^{n\times d_{v}}$, respectively, the attention score is computed as follows:
+    $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V},$ (1)
+[…]
+    $\alpha_{ij}=\text{Softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{j}^{T}}{\sqrt{d_{k}}}\right),$ (2)
+    since $\alpha_{ij}$ signifies the attention weight of the $j$-th input for the $i$-th output.
+With multiple heads, the queries, keys, and values are learned in parallel linear projections, and then concatenated as the output of this layer.
+  • Feed-Forward Neural Network (FFN). A single-layer FFN is utilized to capture non-linear patterns. This layer is represented as:
     $FFN(\bm{x}_{i})=\max(0,\bm{x}_{i}\mathbf{W}_{1}+b_{1})\mathbf{W}_{2}+b_{2},$ (3)
     with weight matrices $\mathbf{W}_{1},\mathbf{W}_{2}$ and bias terms $b_{1},b_{2}$.
-IV-A2 Self-Attention Mechanism
-This module facilitates the model in emphasizing or downplaying certain temporal characteristics.
-The attention weight $\bm{\Lambda}=[\alpha]_{i,j}$ computes the weighted sum with its $(i,j)$-th element defined as:
-$\alpha_{ij}=\text{Softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{j}^{T}}{\sqrt{d_{k}}}\right),$ (2)
-where $\alpha_{ij}$ signifies the attention weight of the $j$-th input for the $i$-th output.
-IV-A3 Decoder Module
-After encoding the time-series data into hidden states, the decoder further refines this information to produce outputs which are conductive for RUL prediction.
+IV-A2 Decoder Module
+After encoding the time-series data into hidden states, the decoder further refines this information to produce outputs that are conducive to RUL prediction.
 Given hidden states $\bm{h}^{\prime}$, the decoder module generates output states $\textbf{O}=\{\bm{o}_{1},\bm{o}_{2},\cdots,\bm{o}_{n}\}$.
+Similar to the encoder, the decoder module is also composed of a number of $L$ identical blocks.
 Each block in the decoder encompasses:
+  • Masked Multi-head Self-attention Mechanism.
+This layer takes the output embedding as the input and is the same as the multi-head self-attention mechanism in the encoder module.
+Masking is applied to make sure that the newly generated output only depends on the historical outputs.
   • Multi-head Cross-Attention Mechanism. Cross-attention takes in the encoder’s outputs and combines them with the decoder’s states. This process is represented as
@@
     where $d_{k}$ represents the dimension of the queries and keys.
-  • Feed-Forward Neural Network. This module extracts patterns in the decoder states:
+Being different from (1), the values of $\mathbf{K}_{e}$ and $\mathbf{V}_{e}$ are derived from the output of the encoder and $\mathbf{Q}_{d}$ is derived from the output embedding of the decoder or the output of the previous block in the decoder.
+  • Feed-forward Neural Network. This module extracts patterns in the decoder states and takes the same form as (3) in the encoder:
     $FFN(\bm{o}_{i})=\max(0,\bm{o}_{i}\mathbf{W}_{3}+b_{3})\mathbf{W}_{4}+b_{4},$
     with weight matrices $\mathbf{W}_{3},\mathbf{W}_{4}$ and biases $b_{3},b_{4}$.
-IV-A4 Prediction Module
-The prediction module utilizes the states $\bm{h}$ to estimate the RUL. This involves one or several dense layers, which can be expressed as
+We note that each attention layer or FFN layer in the encoder and the decoder is accompanied with a residual connection and their output is processed with layer normalization (see Algorithm 2).
+IV-A3 Prediction Module
+The prediction module utilizes the states $\bm{h}$ to estimate the RUL. This involves one or several dense FFN layers, each layer of which can be expressed as
 $\text{RUL}=\sigma(\bm{h}\mathbf{W}+b),$
@@
 return $\mathbf{y}$
-Algorithm 2 A Transformer algorithm for a single machine.
+Algorithm 2 A RUL prediction algorithm for a single machine
+Input: Distributed input sequences $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{K}$ from $K$ machines
+Output: Aggregated output sequences $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{K}$
+1
+2Initialize a central model with positional encoding and input embeddings;
+3for each training round do
+       /* Local Training on Each Machine */
+4       for each machine $k=1$ to $K$ do
+5             Receive model parameters from the central model;
+6             Apply Algorithm 2 with the given number of layers on local sequence $\mathbf{x}_{k}$;
+7             Compute local updates and send them back to the central model;
+8
+9       end for
+10
+      /* Aggregation at the Central Model */
+11       Aggregate local updates from all machines;
+12       Update central model parameters;
+13
+14 end for
+15
+/* Output Generation */
+16 for each machine $k=1$ to $K$ do
+17       Send final model parameters to machine $k$;
+18       Generate output sequence $\mathbf{y}_{k}$ using the final model on local input $\mathbf{x}_{k}$;
+19
+20 end for
+21
+return $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{K}$.
+Algorithm 3 A federated RUL prediction algorithm for multiple machines
 IV-B Case 2: A Federated RUL Prediction Module for Multi-Machines
-The multi-machine case is featured by a federated model, which dispatches an individual training model for similar machines.
-The local model takes as input the local time series data $\mathbf{X}$ and outputs an estimation of the RUL $R(t)$.
+The multi-machine case is featured by a federated model, which dispatches an individual training model for $K$ similar machines.
+A local model $k$ takes as input the local time series data $\mathbf{X}_{k}$ and outputs an estimation of the RUL $R_{k}(t)$ regarding the machine it concerns.
+With sensors deployed in the same way, different local models observe time-series inputs of the same dimension and share the same prediction label set.
 By combining the strengths of multiple models extracted from different machines, the approach improves the reliability and accuracy of RUL predictions, enabling industries to make informed decisions regarding maintenance schedules and machinery operations.
-The averaged model for RUL prediction leverages an ensemble of predictive models to provide a more precise estimate of RUL, enhancing the efficiency and reliability of predictive maintenance in industries.
+The global model for RUL prediction leverages an ensemble of predictive models to provide a more precise estimate of RUL, enhancing the efficiency and reliability of predictive maintenance in industries.
 As shown in the Algorithm 3, given a set of time series data $\mathbf{X}$, the ensemble model aims to predict the RUL $R(t)$ of multiple machines.
@@
 This enhanced prediction quality allows industries to make precise maintenance schedules and optimize machine operations.
-The averaged model for RUL prediction can be represented as:
-$R(t)=\frac{1}{m}\sum_{i=1}^{m}f_{i}(\mathbf{X}),$ (5)
-where $m$ is the number of models in the ensemble, and $f_{i}$ represents the individual predictive model.
+The global model for RUL prediction can be represented as
+$R(t)=\frac{1}{K}\sum_{k=1}^{K}f_{k}(\mathbf{X}),$ (6)
+where $K$ is the number of models in the ensemble, and $f_{k}$ represents the individual predictive model.
 This averaging method capitalizes on the collective strength of diverse predictive algorithms.
 Based on the predictive RULs, the infrastructure can conduct equipment maintenance or replacement to ensure uninterrupted and reliable functionality of the machines.
-Input: Distributed input sequences $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{K}$ from $K$ machines
-Output: Aggregated output sequences $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{K}$
-1 Initialize a central model with positional encoding and input embeddings; for each training round do
-       /* Local Training on Each Machine */
-2       for each machine $k=1$ to $K$ do
-3             Receive model parameters from the central model;
-4             Apply Algorithm 2 with the given number of layers on local sequence $\mathbf{x}_{k}$;
-5             Compute local updates and send them back to the central model;
-6
-7       end for
-      /* Aggregation at the Central Model */
-8       Aggregate local updates from all machines;
-9       Update central model parameters;
-10
-11 end for
-/* Output Generation */
-12 for each machine $k=1$ to $K$ do
-13       Send final model parameters to machine $k$;
-14       Generate output sequence $\mathbf{y}_{k}$ using the final model on local input $\mathbf{x}_{k}$;
-15
-16 end for
-return $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{K}$
-Algorithm 3 A Federated Transformer Algorithm for Multiple Machines
 The distinction between Case 1 and Case 2 lies primarily in their scale and data handling methods.
@@
 This method aggregates data from various machines, synthesizing diverse inputs for a broader analysis.
-It suits for a industry which contains multiple similar production lines or machines.
-Additionally, Case 2 could apply Case 1’s algorithm to each machine and then focus on multi-machine data aggregation.
-IV-C Maintenance Action Recommendation Module
+It suits an industry that contains multiple similar production lines or machines.
+Additionally, Case 2 can suitably apply Case 1’s algorithm to each machine and then focus on multi-machine data aggregation.
+Input: Exploration rate $\epsilon$ (for DQN, SAC), learning rate $\alpha$, discount factor $\gamma$ and policy update frequency (for PPO, SAC).
+Output: Optimal action $a$ for maintenance
+1
+/* Initialization */
+2 Initialize the policy network and optionally value/critic network(s) with the input hyperparameters;
+3 Initialize target network(s) if applicable (e.g., for DQN, SAC);
+4 Initialize experience replay buffer if applicable;
+5
+6for each episode do
+7       Initialize the starting state $\bm{s}$;
+8       while $\bm{s}$ is not a terminal state do
+             /* Action Selection */
+9             if using DQN or SAC then
+10                   Select action $a$ using an $\epsilon$-greedy policy or according to the current policy and exploration strategy;
+11
+12             else if using PPO then
+13                   Select action $a$ according to the current policy;
+14
+15             end if
+16            Execute action $a$, observe reward $r$ and successor state $\bm{s^{\prime}}$;
+17
+            /* Policy and (or) Value Network Update */
+18             if using DQN then
+19                   Store transition $(\bm{s},a,r,\bm{s^{\prime}})$ in replay buffer;
+20                   Sample mini-batch and update the network(s) using TD error;
+21                   Periodically update target network;
+22
+23             else if using PPO then
+24                   Accumulate data for policy update;
+25                   Update policy using PPO;
+26
+27             else if using SAC then
+28                   Store transition in replay buffer;
+29                   Sample mini-batch and update policy and value networks using SAC update rules;
+30                   Adjust temperature parameter for entropy maximization;
+31
+32             end if
+33
+            /* Exploration Rate Adjustment */
+34             Adjust the exploration rate $\epsilon$ if applicable;
+35             Update current state $\bm{s}\leftarrow\bm{s^{\prime}}$;
+36
+37       end while
+      /* Policy Evaluation */
+38       Periodically evaluate the performance of the current policy;
+39
+40 end for
+return Optimal maintenance action $a$
+Algorithm 4 A DRL-based algorithm for optimal maintenance action recommendation
+IV-C Deep Reinforcement Learning for Maintenance Action Recommendation
 Upon obtaining the RUL predictions from the previous module, the maintenance action recommendation module initiates.
@@
 For instance, should the RUL prediction for a component fall below a designed limit (e.g., 10 hours), an immediate maintenance notification is dispatched.
-On the other hand, if the RUL surpasses the limit, periodic inspections is considered sufficient.
+On the other hand, if the RUL surpasses the limit, periodic inspections are considered sufficient.
 Strategy-Learning Agents.
@@
 Inputting the present RUL prediction enables the agent to recommend the most probable efficient maintenance actions, grounded in historical insights.
-The appropriate integration of the RUL predictions with the maintenance action recommendation module ensures that maintenance schedules is efficient and data-driven, ultimately leading to improved system reliability, longer equipment lifespan, and reduced operational costs.
+The appropriate integration of the RUL predictions with the maintenance action recommendation module ensures that maintenance schedules are efficient and data-driven, ultimately leading to improved system reliability, longer equipment lifespan, and reduced operational costs.
 Thus, the predicted RUL, indicative of a system’s operational state, is assimilated into the state representation within a Markov Decision Process (MDP)-based framework.
-The subsequent discussions elaborate the maintenance strategy-learning procedure under the MDP framework, which includes following main components:
+The subsequent discussions elaborate on the maintenance strategy-learning procedure under the MDP framework, which includes the following components:
 Environment State Space.
-As established, the system can be in one of $n$ distinct degradation states, with $\mathbf{S}^{t}=[S_{1}^{t},S_{2}^{t},\ldots,S_{n}^{t}]$ consisting all possible $n$ states that the system can take at time $t$.
+As established, the system can be in one of $n$ distinct degradation states, with $\mathbf{S}^{t}=\{S_{1}^{t},S_{2}^{t},\ldots,S_{n}^{t}\}$ consisting all possible $n$ states that the system can take at time $t$.
 Action Space. If we denote the number of possible maintenance actions by $m$, the action space $\mathcal{A}$ can be represented as $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{m}\}$.
@@
 Policy.
-A policy $\pi$ is a mapping from states to actions.
+A policy $\pi$ is a (probabilistic) mapping from states to actions.
 An optimal policy $\pi^{*}$ maximizes the expected cumulative discounted reward from any starting state.
@@
 Optimal Policies under Bellman Equation. The value of a state under an optimal policy $\pi^{*}$ obeys the following recursive Bellman optimality equation
-$V^{*}(s_{i})=\max_{a\in A}\left[R(s_{i},a)+\gamma\sum_{j=1}^{n}P_{i,j}^{a}V^{*}(s_{j})\right].$
-On top of this MDP-based policy derivation model, we are able to develop a DRL-based plug-and-play module for maintenance action recommendation using neural network-based function/policy approximation, by accommodating a number of state-of-the-art DRL algorithms.
-Algorithm 4 shows the details of leveraging different DRL algorithms, including DQN, SAC, PPO, for selecting optimal maintenance action $a$ through evaluating the performance of each policy.
-Details of hyperparameters setting and how to apply the algorithm to the real-world case are explained in Section V-C4.
-Input: Exploration rate $\epsilon$ (for DQN, SAC), learning rate $\alpha$, discount factor $\gamma$ and policy update frequency (for PPO, SAC)
-Output: Optimal action $a$ for maintenance
-/* Initialization */
-1 Initialize the policy network and optionally value/critic network(s) with the input hyperparameters;
-2 Initialize target network(s) if applicable (e.g., for DQN, SAC);
-3 Initialize replay buffer if applicable;
-4 for each episode do
-5       Initialize the starting state $\bm{s}$;
-6       while $\bm{s}$ is not a terminal state do
-             /* Action Selection */
-7             if using DQN or SAC then
-8                   Select action $a$ using an $\epsilon$-greedy policy or according to the current policy and exploration strategy;
-9
-10             else if using PPO then
-11                   Select action $a$ according to the current policy;
-12
-13             end if
-14            Execute action $a$, observe reward $r$ and successor state $\bm{s^{\prime}}$;
-             /* Policy and (or) Value Network Update */
-15             if using DQN then
-16                   Store transition $(\bm{s},a,r,\bm{s^{\prime}})$ in replay buffer;
-17                   Sample mini-batch and update the network(s) using TD error;
-18                   Periodically update target network;
-19
-20             else if using PPO then
-21                   Accumulate data for policy update;
-22                   Update policy using PPO;
-23
-24             else if using SAC then
-25                   Store transition in replay buffer;
-26                   Sample mini-batch and update policy and value networks using SAC update rules;
-27                   Adjust temperature parameter for entropy maximization;
-28
-29             end if
-            /* Exploration Rate Adjustment */
-30             Adjust the exploration rate $\epsilon$ if applicable;
-31             Update current state $\bm{s}\leftarrow\bm{s^{\prime}}$;
-32
-33       end while
-      /* Policy Evaluation */
-34       Periodically evaluate the performance of the current policy;
-35
-36 end for
-return Optimal maintenance action $a$
-Algorithm 4 A Deep Reinforcement Learning Algorithm for Maintenance Action Recommendation
-V Case Study
-In this section, we perform a case study on aircraft engine maintenance using NASA’s C-MAPASS dataset [28] to evaluate our proposed TranDRL framework.
-V-A Dataset
-The NASA’s C-MAPSS dataset [28] contains four subsets denoted as FD001, FD002, FD003, and FD004, each having a training and test set, reflecting distinct operating and fault modes.
-The training data comprises inputs from three operating conditions and 21 sensors’ measurements.
-For training, every engine data point serves as a sample paired with its RUL.
-In the experiment, we applied a piecewise linear degradation model [29] for RUL determinations.
-The last cycle data point for each engine generally stands as the sample, with actual RUL values included in the dataset.
-Figure 2: An Overview of the Temporal Fusion Transformer.
-V-B Remaining Useful Life Prediction
-The framework given in Algorithm 3 makes it suitable to incorporate the Transformer-based time-series event forecasting networks with other feature extraction and post-processing modules.
-In the case study, we adopt the Temporal Fusion Transformer (TFT) [30] in Algorithm 3 to utilizes the advantages of both recurrent neural networks and Transformers [31] into a unified encoder-decoder architecture.
-Compared to traditional time series forecasting models, e.g., Long Short-Term Memory networks (LSTMs) [32], TFT offers enhanced interpretability and the capability to deal with multiple prediction horizons.
-The structure of TFT is shown in the Figure 2 and the includes following components:
-Encoder. The encoder transforms the raw time-series data into a high-dimensional space using convolutional layers and self-attention mechanisms, thereby learning rich representations of the input.
-Temporal Attention. This specialized module let the model focus on different time-steps, facilitating the capture of dynamic temporal patterns and long-range dependencies.
-Decoder. Tailored for multi-horizon prediction, the decoder employs an autoregressive model that recursively uses its previous forecasts as inputs for forecasting subsequent steps.
-Static Inputs. Apart from time-dependent features, TFT can also integrate static inputs to provide a comprehensive context for forecasting.
-Interpretable Components. To ensure transparency, TFT integrates mechanisms like gating and linear autoregressive decoders that allow for human-readable prediction interpretation.
-Evaluation Results.
-In our comparisons, Figures 3 to 6 demonstrate the TFT algorithm’s superior performance over NBeats [33] and DeepAR [34] in predicting the RUL for a individual machine, as evidenced by consistently higher RUL values.
-This suggests TFT’s enhanced capability in capturing complex, time-dependent patterns in sensor data, highlighting its robustness across various machine behaviors.
-In contrast, Figure 7 shows that while federated learning models generate reasonably good average RUL values across multiple machines, they may not capture the intricate details as effectively as single-machine models.
-However, Figures 8 to 11 show a significant bias when models trained on a single machine are applied to other machines, underscoring the challenge of model generalizability in diverse operational environments.
-This analysis illustrates the trade-off between the granularity and specificity of single-machine models and the broader applicability but potentially lower precision of federated learning approaches, highlighting the need for context-aware modeling strategies in predictive maintenance.
-Figure 3: The RUL prediction result for machine FD001.
-Figure 4: The RUL prediction result for machine FD002.
-Figure 5: The RUL prediction result for machine FD003.
-Figure 6: The RUL prediction result for machine FD004.
-Figure 7: Comparison of RUL prediction results of four machines using the averaged model.
-Figure 8: Comparison of RUL prediction results of four machines using the model trained from machine FD001.
-Figure 9: Comparison of RUL prediction results of four machines using the model trained from machine FD002.
-Figure 10: Comparison of RUL prediction results of four machines using the model trained from machine FD003.
-Figure 11: Comparison of RUL prediction results of four machines using the model trained from machine FD004.
-V-C Optimal Maintenance Actions Recommendation
-In the RUL prediction stage, experiment results show that our proposed framework succeeds in predicting RULs using four datasets from NASA C-MAPSS dataset [28].
+$V^{*}(s_{i})=\max_{a\in A}\left[R(s_{i},a)+\gamma\sum_{j=1}^{n}P_{i,j}^{a}V^{*}(s_{j})\right].$ (7)
+On top of this MDP-based policy derivation model, we can develop a DRL-based plug-and-play module for maintenance action recommendation using neural network-based function/policy approximation, by accommodating many state-of-the-art DRL algorithms.
+Algorithm 4 shows details of leveraging different DRL algorithms, including DQN, SAC, and PPO (see our later discussion for the details of each algorithm), for selecting optimal maintenance action $a$ through evaluating the performance of each policy.
+In the RUL prediction stage, experiment results in Section V show that our proposed framework succeeds in predicting RULs using four sub-datasets in the NASA C-MAPSS open dataset [26].
 However, a gap still exists between the prediction results and the maintenance action decisions needed by engineers.
-To make action recommendation by weighing the cost and benefits, we leverage DRL algorithms as given in Algorithm 4, which helps to find the best trade-off.
-V-C1 States and Actions
-The state space is generated from PCA features of the RUL resulted from the prediction outcomes of the NASA’s C-MAPSS Aircraft Engine Simulator dataset.
-These PCA features provide a compact and important part of the machine’s current health condition.
+To make action recommendations by weighing the cost and benefits, we leverage DRL algorithms as given in Algorithm 4, which helps to find the best trade-off.
+IV-C1 States and Actions
+The state space is generated from PCA features of the RUL resulting from the prediction outcomes of NASA’s C-MAPSS Aircraft Engine Simulator dataset.
+These PCA features provide a compact and important part of the current health condition of the machine.
 The action space contains three actions:
@@
 The DRL agent aims to navigate this state-action space to minimize long-term operational costs by selecting the optimal maintenance actions based on each state’s representation of machine health.
-V-C2 Probabilities Model
-In our study, the transition and action cost matrices are empirically calibrated using the NASA C-MAPSS dataset.
+IV-C2 Probabilities Model
+For the purpose of evaluating the performance of our proposed algorithm, the transition and action cost matrices are empirically calibrated using the NASA C-MAPSS dataset to generate the controllable ground truth.
 Specifically, each entry in the state transition matrix $p(s_{t+1}|s_{t},a_{t})$ reflects the probability of transitioning from state $s_{t}$ to $s_{t+1}$ when action $a_{t}$ is taken, and these probabilities are derived through statistical analyses of state sequences and their corresponding RUL values in the C-MAPSS dataset.
@@
 In this MDP framework, the transition matrix has a dimension $(10,3,10)$, representing transitions between $10$ discretized states given $3$ actions, all closely reflecting scenarios that commercial aero-engines might encounter.
-V-C3 Reward Model
-The objective of maintenance actions recommend algorithms is not only to maximize the RUL but also to minimize the operational costs and downtime.
+IV-C3 Reward Model
+The objective of maintenance actions recommends algorithms not only to maximize the RUL but also to minimize operational costs and downtime.
 To fulfill these multi-objective requirements, we propose a reward function that combines these elements in a weighted manner.
-The reward function $R(S_{t},a_{t},S_{t+1})$ is formulated as:
+The reward function $R(s_{t},a_{t},s_{t+1})$ is formulated as
 $R(s_{t},a_{t},s_{t+1})$ $=\alpha\times\text{RUL}_{\text{gain}}(s_{t},a_{t})~{}$
@@
 A higher RUL gain is desirable as it suggests an extension in the machine’s normal working time.
-  • $\text{Cost}(a_{t})$ denotes the operational or maintenance cost incurred by taking action $a_{t}$. This cost is modeled based on experiences.
+  • $\text{Cost}(a_{t})$ denotes the operational or maintenance cost incurred by taking action $a_{t}$. This cost is modeled based on experience.
   • $\text{DownTime}(a_{t})$ signifies the machine downtime resulting from action $a_{t}$. Downtime is critical in operational settings, which impacts productivity or service levels.
@@
 By fine-tuning these hyperparameters, the model can be customized to focus on extending the RUL, minimizing costs, or reducing downtime.
-V-C4 Hyperparamters
-For each DRL algorithm embedded in Algorithm 4, we carefully choose its corresponding hyperparameters based on empirical experiments. Details for each DRL algorithm are given as follows.
+IV-D Integrate with Human Feedback for Maintenance Action Recommendation
+Incorporating human expertise into the reinforcement learning process helps to resolve some of the challenges encountered with traditional RL methods.
+Especially, in scenarios where real-world rewards are hard to quantify or can be misleading, human feedback provides a way to guide the learning process more effectively.
+Thus, as machines and their associated systems become increasingly complex, conditions that need to be considered may take a long time for Algorithm 4.
+Human experts, with their working experiences and domain knowledge to find complex patterns, can provide valuable insights that may not be readily available through raw data.
+To minimize labor costs, we assume that one expert will be involved during the human feedback process.
+By integrating human feedback into the RL process, the framework can:
+  • Augment the learning process with domain-specific knowledge.
+  • Address the cold start problem where initial exploration might be risky or expensive.
+  • Improve learning efficiency by reducing the exploration of sub-optimal actions.
+The above goals are achieved by incorporating the following information into our proposed framework in Algorithm 5:
+  1. 1.
+    Human Opinions: Collect human predictions of the desired behavior.
+Experienced engineers can provide opinions on maintenance decisions based on various machine states obtained from the C-MAPSS dataset.
+  2. 2.
+    Feedback on Agent’s Actions: After the agent makes a decision, a human can provide feedback on the chosen action, guiding the agent on the appropriate actions.
+Then, we introduce a term in the reward function that captures human feedback
+$R_{\text{human}}(a_{t})=\begin{cases}r_{\text{positive}},&\text{if human feedback is positive},\\ r_{\text{negative}},&\text{if human feedback is negative},\\ 0,&\text{otherwise},\end{cases}$ (9)
+based on which, the modified reward function from (IV-C3) for DRL becomes
+$R(S_{t},a_{t},S_{t+1})$ $=\alpha\times\text{RUL}_{\text{gain}}(S_{t},a_{t})-\beta\times\text{Cost}(a_{t})-$
+$\gamma\times\text{DownTime}(a_{t})+\delta\times R_{\text{human}}(a_{t}),$
+where $\delta$ is a hyperparameter weighing the importance of human feedback.
+Input: Observation $\bm{s}$, Exploration rate $\epsilon$, Learning rate $\alpha$, Discount factor $\gamma$, Human feedback 00.
+Output: Optimal action $a$ for maintenance
+1
+2Initialize the neural network weights;
+3 Initialize a mechanism to incorporate human feedback (e.g., reward shaping, action recommendation);
+4
+5for each episode do
+       /* Incorporate Human Demonstrations */
+6       If available, utilize human demonstrations to set initial policy;
+7
+8      Initialize the starting state $\bm{s}$;
+9       while $s$ is not a terminal state do
+             /* Action Selection with Human Feedback */
+10             Select action $a$ using an $\epsilon$-greedy policy or based on human recommendations;
+11             Execute action $a$, observe reward $r$ and successor state $\bm{s^{\prime}}$;
+12
+            /* Adjust Reward Based on Human Feedback */
+13             Incorporate human feedback into the reward:
+$R_{\text{human}}(a_{t})=\begin{cases}r_{\text{positive}},&\text{if HF is positive},\\ r_{\text{negative}},&\text{if HF is negative},\\ 0,&\text{otherwise}.\end{cases}$
+Adjust reward $r$ based on $R_{\text{human}}(a_{t})$;
+14
+            /* Network Update */
+15             Calculate TD error $\delta=r+\gamma\max_{a^{\prime}}Q(\bm{s^{\prime}},a^{\prime})-Q(\bm{s},a)$;
+16             Update neural network weights using $\delta$ and learning rate $\alpha$;
+17
+            /* Adjustment of Exploration and State */
+18             Adjust the exploration rate $\epsilon$;
+19             Update current state $\bm{s}\leftarrow\bm{s^{\prime}}$;
+20
+21       end while
+      /* Incorporate Feedback into Learning Process */
+22       Periodically adjust the learning process based on accumulated human feedback;
+23
+24 end for
+return Optimal maintenance action $a$.
+Algorithm 5 A RLHF-based algorithm for optimal maintenance action recommendation
+V Case Study
+In this section, we present a case study of our proposed TranDRL framework.
+V-A Data
+NASA’s C-MAPSS dataset [26] contains four subsets denoted as FD001, FD002, FD003, and FD004, each with a training and test set, reflecting distinct operating and fault modes.
+The training data comprise inputs from $3$ operating conditions and $21$ sensor measurements across four engines of the same type.
+While the datasets are i.i.d, each engine has its unique operational settings and sensor readings.
+By leveraging the C-MAPSS dataset, we can gain insights into how IoT systems might perform in real-world scenarios, particularly those involving complex machinery such as aircraft engines.
+The dataset facilitates data collection and monitoring, mimicking real-world engine performance, which represents the general working scenario for IoT-based sensing and decision making in different industrial sectors.
+Therefore, it provides a convenient platform for the development and testing of predictive maintenance and anomaly detection algorithms by offering simulated engine degradation and fault data.
+V-B Experiment Setup
+In the experiment, we applied a piecewise linear degradation model [27] for RUL determinations.
+Each engine’s last cycle data point generally stands as the sample, with actual RUL values included in the dataset.
+Environment Settings. We utilized an Intel Core i9-14900K CPU processor, 128GB DDR5 RAM, and a PNY RTX 4090 GPU card. The system ran on Ubuntu 22.04 LTS with Python 3.9 and PyTorch 2.0.1 for tasks.
+Hyperparamters. We carefully choose its corresponding hyperparameters based on empirical experiments.
+Transformers.
+The TFT model employs a learning rate of 0.03, hidden layers of size 16, a single attention head, a dropout rate of 0.1, a hidden continuous size of 8, and Symmetric Mean Absolute Percentage Error (SMAPE) as the loss function, focusing on the target variable RUL.
+We also introduce two baseline algorithms for prediction performance comparison.
+For the first reference algorithm, the N-Beats model [28], also set with a learning rate of 0.03, uses an encoder length of 120, a prediction length of 60, a batch size of 128, implement early stopping at 10 epochs, and uses gradient clipping at 0.1.
+For another reference algorithm, the DeepAR model [29], parameters include a lower learning rate of 0.003, encoder and decoder lengths of 120 and 60 respectively, the same batch size of 128, target normalization via GroupNormalizer, training up to 10 epochs with early stopping after 4, and weight decay at 0.01.
+The negative binomial distribution loss is employed.
 Deep Q-network (DQN).
@@
 The learning rate is set to $10^{-4}$, the discount factor is set to $0.99$, and the replay buffer size is set to $10^{6}$ experiences.
-We use a batch size of $512$ and update the target network every $50$ steps.
+We use a batch size of $256$ and update the target network every $50$ step.
 The exploration-exploitation trade-off is controlled by an $\epsilon$ value that started at $1.0$ and decayed over time to a final value of $0.01$.
@@
 A clip range of $0.2$ is used to limit the updates to the policy, and we apply a value function coefficient of $0.5$ to balance the importance of the policy and the value function.
-Additionally, we set the entropy coefficient to $0.01$ to encourage exploration and a batch size as $64$.
-Then, we optimize the policy every 10 epochs and use the Adam optimizer with a learning rate of $0.0001$.
-Finally, GAE parameter is set to $0.95$ to estimate the advantage function.
+Additionally, we set the entropy coefficient to $0.01$ to encourage exploration and a batch size of $256$.
+Then, we optimize the policy every 10 epochs and use the Adam optimizer with a learning rate of $10^{-4}$.
+Finally, the GAE parameter is set to $0.95$ to estimate the advantage function.
 Soft Actor-Critic (SAC).
@@
 SAC aims to maximize an objective function that balances both the expected return and the entropy of the policy, making it more robust and efficient.
-The learning rate for the actor and critic networks is set to $0.0003$, and we use a discount factor of $0.99$.
-Then, we apply a temperature parameter alpha of $0.2$ to balance the entropy term in the objective function.
+The learning rate for the actor and critic networks is set to $10^{-4}$, and we use a discount factor of $0.99$.
+Then, we apply a temperature parameter $\alpha$ of $0.2$ to balance the entropy term in the objective function.
 The target entropy is set to $-2.0$, and we use a replay buffer of $10^{6}$ experiences with a batch size of $256$.
 We also apply a soft update of the target networks with a rate of $0.005$, and we initialize the log standard deviation of the policy with a value of $-2$.
-Finally, we train the algorithm for $10^{5}$ time steps and use the Adam optimizer with a learning rate of $0.0001$.
-V-D Results and Discussion
-Figure 12: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
-Figure 13: The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
-We compare the performance of DQN, PPO, and SAC on our designed task in Figures 1213.
-The evaluation metric is the total rewards obtained by the agent during a fixed number of episodes.
-The experimental results show that PPO outperforms both DQN and SAC, achieving higher total rewards on the task.
-The average total rewards obtained by PPO are $55,000$, while DQN and SAC achieve average total rewards of $50,000$ and $45,000$, respectively.
+Finally, we train the algorithm for $10^{5}$ time steps and use the Adam optimizer with a learning rate of $10^{-4}$.
+(a) The RUL prediction results for machine FD001.
+(b) The RUL prediction results for machine FD002.
+(c) The RUL prediction results for machine FD003.
+(d) The RUL prediction results for machine FD004.
+Figure 2: RUL prediction results for four machines.
+Baseline.
+The ground truth is obtained from run-to-failure data, where the machine is operated until failure occurs.
+The run-to-failure data consists of time-series measurements of various sensor data recorded during operation until failure.
+The ground truth RUL values are calculated by subtracting the current cycle $t$ from the cycle $T_{f}$ at which failure occurred, expressed mathematically as $\text{RUL}(t)=T_{f}-t$.
+For baseline methods, we utilize NBeats [28] and DeepAR [29] to compare against our proposed method, aiming to verify whether it outperforms the baselines.
+V-C Results and Analysis
+Figures 2(a)-2(d) demonstrate the superior performance of our proposed algorithm (Algorithm 2 as TFT) over NBeats and DeepAR in predicting the RUL for an individual machine, as evidenced by consistently higher RUL values.
+This suggests TFT’s enhanced capability in capturing complex, time-dependent patterns in high dimensional sensor data, highlighting its robustness across various machine behaviors.
+(a) Comparison of RUL prediction results of four machines using the model trained from machine FD001.
+(b) Comparison of RUL prediction results of four machines using the model trained from machine FD002.
+(c) Comparison of RUL prediction results of four machines using the model trained from machine FD003.
+(d) Comparison of RUL prediction results of four machines using the model trained from machine FD004.
+(e) Comparison of RUL prediction results of four machines using the global model.
+Figure 3: Comparison of RUL prediction results of four machines using the model.
+In contrast, Figure 3(e) shows that while the federated learning models (Algorithm 3) generate reasonably good average RUL values across multiple machines, they may not capture the intricate details as effectively as single-machine models.
+However, Figures 3(a) to 3(d) show a significant bias when models trained on a single machine are applied to other machines, underscoring the challenge of model generalizability in diverse operational environments.
+This analysis illustrates the trade-off between the granularity and specificity of single-machine models and the broader applicability but potentially lower precision of federated learning approaches, highlighting the need for context-aware modeling strategies in predictive maintenance.
+(a) The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
+(b) The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
+Figure 4: Results of Using Deep Reinforcement Learning for Optimizing Maintenance Strategies.
+(a) The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
+(b) The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
+Figure 5: Results for Using Reinforcement Learning with Human Feedback for Optimizing Maintenance Strategies.
+We compare the performance of DQN, PPO, and SAC in the framework of TranDRL (Algorithm 4) in our designed maintenance task in Figures 4(a) and 4(b).
+The evaluation metric is the total reward obtained by the agent during a fixed number of episodes.
+The experimental results show that PPO outperforms both DQN and SAC, achieving a higher total reward on the task.
+The average total reward obtained by PPO is $55,000$, while DQN and SAC achieve an average total reward of $50,000$ and $45,000$, respectively.
 However, DQN converges faster than PPO and SAC.
@@
 On the other hand, if the action space is continuous or the observations are high-dimensional, PPO and SAC may be more suitable and converge faster.
-Overall, these experimental results demonstrate the importance of selecting the appropriate DRL algorithm and hyperparameters for a specific task to achieve the optimal performance.
-Figure 14: Differences between RL and RLHF.
-V-E Reinforcement Learning with Human Feedback
-Incorporating human expertise into the reinforcement learning process helps to resolve some of the challenges encountered with traditional RL methods.
-Especially, in scenarios where real-world rewards are hard to quantify or can be misleading, human feedback provides a way to guide the learning process more effectively.
-The comparison between RL and RLHF is presented in Fig.
-14.
-As machines and their associated systems become increasingly complex, the conditions that need to be considered maybe take a long time for Algorithm 5.
-Human experts, with their working experiences and domain knowledge to find complex patterns, can provide valuable insights that may not be readily available through raw data.
-By integrating human feedback into the RL process, the framework can:
-  • Augment the learning process with domain-specific knowledge.
-  • Address the cold start problem where initial exploration might be risky or expensive.
-  • Improve learning efficiency by reducing the exploration of sub-optimal actions.
-Thus, we further incorporate the following information into our proposed framework in Algorithm 5:
-  1. 1.
-    Human Opinions: Collect human predictions of the desired behavior. Experienced engineers can provide opinions on maintenance decisions based on various machine states from the C-MAPSS dataset.
-  2. 2.
-    Feedback on Agent’s Actions: After the agent makes a decision, a human can provide feedback on the chosen action, guiding the agent on the appropriateness of its actions.
-Then, we introduce a term in the reward function that captures human feedback
-$R_{\text{human}}(a_{t})=\begin{cases}r_{\text{positive}},&\text{if human feedback is positive},\\ r_{\text{negative}},&\text{if human feedback is negative},\\ 0,&\text{otherwise},\end{cases}$ (7)
-based on which, the modified reward function for DRL is
-$R(S_{t},a_{t},S_{t+1})$ $=\alpha\times\text{RUL}_{\text{gain}}(S_{t},a_{t})-\beta\times\text{Cost}(a_{t})-$
-$\gamma\times\text{DownTime}(a_{t})+\delta\times R_{\text{human}}(a_{t}),$
-where $\delta$ is a hyperparameter weighing the importance of human feedback.
-Input: Observation $\bm{s}$, Exploration rate $\epsilon$, Learning rate $\alpha$, Discount factor $\gamma$, Human feedback 00.
-Output: Optimal action $a$ for maintenance
-1 Initialize the neural network weights;
-2 Initialize a mechanism to incorporate human feedback (e.g., reward shaping, action recommendation);
-3 for each episode do
-       /* Incorporate Human Demonstrations */
-4       If available, utilize human demonstrations to set initial policy;
-5       Initialize the starting state $\bm{s}$;
-6       while $s$ is not a terminal state do
-             /* Action Selection with Human Feedback */
-7             Select action $a$ using an $\epsilon$-greedy policy or based on human recommendations;
-8             Execute action $a$, observe reward $r$ and successor state $\bm{s^{\prime}}$;
-             /* Adjust Reward Based on Human Feedback */
-9             Incorporate human feedback into the reward:
-$R_{\text{human}}(a_{t})=\begin{cases}r_{\text{positive}},&\text{if HF is positive},\\ r_{\text{negative}},&\text{if HF is negative},\\ 0,&\text{otherwise}.\end{cases}$
-Adjust reward $r$ based on $R_{\text{human}}(a_{t})$;
-             /* Network Update */
-10             Calculate TD error $\delta=r+\gamma\max_{a^{\prime}}Q(\bm{s^{\prime}},a^{\prime})-Q(\bm{s},a)$;
-11             Update neural network weights using $\delta$ and learning rate $\alpha$;
-             /* Adjustment of Exploration and State */
-12             Adjust the exploration rate $\epsilon$;
-13             Update current state $\bm{s}\leftarrow\bm{s^{\prime}}$;
-14
-15       end while
-      /* Incorporate Feedback into Learning Process */
-16       Periodically adjust the learning process based on accumulated human feedback;
-17
-18 end for
-return Optimal maintenance action $a$.
-Algorithm 5 The RLHF Algorithm for Maintenance Actions Selection
-V-F Results for RLHF
-Figure 15: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
-Figure 16: The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
-Figure 15 and Figure 16 reflect the changes after adding on the human feedback in the reward function.
+Overall, these experimental results demonstrate the importance of selecting the appropriate DRL algorithm and hyperparameters for a specific task to achieve optimal performance.
+Figure 5(a) and Figure 5(b) reflect the changes after adding on the human feedback in the reward function (Algorithm 5).
 The results show that the speed of convergence does not improve too much, but the process becomes more stable and variances minimized.
@@
 In the future, we will conduct experiments on more complex real-world scenarios to confirm the effectiveness of the RLHF algorithm.
-VI Conclusions
-In this paper, we present a comprehensive prescriptive maintenance framework that integrates real-time RUL estimations with DRL algorithms for dynamic action recommendations.
-The framework starts by employing the Transformer network to estimate the RUL distribution of engines, offering a robust and adaptive data-driven approach to understanding machine conditions in real-time.
-As machine operates, our model continues to dynamically updating RUL estimates, providing an ever-evolving status of the equipment’s condition.
-We rigorously compare the performance of several state-of-the-art time-series forecasting models, including TFT, NBeats, and DeepAR, to validate the accuracy of our RUL estimates.
-Our findings indicate that the Transformer-based algorithm delivers superior predictive performance, laying a solid foundation for the subsequent decision-making process.
-For DRL-based maintenance action recommendation, we demonstrate the efficiency of SAC, DQN, and PPO.
-Each of these algorithms has its strengths and weaknesses, but they all are adapted to working seamlessly with the dynamically updated RUL estimates.
-Future work may involve integrating more complex cost models, incorporating uncertainty estimates, and testing the framework in a multi-agent setting or across different types of machines.
-In conclusion, our work signifies a step forward in the request for more intelligent, adaptive, and cost-effective prescriptive maintenance systems.
-By leveraging cutting-edge algorithms for both RUL estimation and action recommendation, we offer a solution that dynamically adapts to the operational realities, thereby enhancing both economic and operational efficiencies.
+VI Conclusion and Future Work
+In this paper, we present a prescriptive maintenance framework (TranDRL) that integrates real-time RUL estimations with DRL algorithms for dynamic action recommendations.
+The framework starts by employing the Transformer network to estimate the RUL distribution of machines (e.g., engines), offering a robust and adaptive data-driven approach to understanding machine conditions in real-time.
+As the machine operates, our model continues to dynamically update RUL estimates, providing an ever-evolving status of the equipment’s condition.
+We rigorously compare the performance of several state-of-the-art time-series forecasting models, including NBeats and DeepAR, with our proposed method to validate the accuracy of our RUL estimates.
+For the framework of DRL-based maintenance action recommendations, we demonstrate the efficiency of SAC, DQN, and PPO with human feedback.
+Each of these algorithms has its strengths and weaknesses, but they are all adapted to working seamlessly with the dynamically updated RUL estimates.
+The limitations of our framework include its inability to process image-based data and the lack of metrics to evaluate the accuracy of human experts’ contributions.
+These shortcomings may compromise the performance of the framework.
+Future work will focus on refining the framework, such as introducing new neural network designs, to address these limitations.